### v2

# January 10, 2025

Abstract.
In this paper we justify Koiter’s model for linearly elastic generalised membrane shells of the first kind subjected to remaining confined in a prescribed half-space. After showing that the confinement condition considered in this paper is in general stronger, under the validity of the Kirchhoff-Love assumptions, than the classical Signorini condition, we formulate the corresponding obstacle problem for a three-dimensional linearly elastic generalised membrane shell of the first kind, and we conduct a rigorous asymptotic analysis as the thickness approaches to zero on the unique solution for one such model. We show that the solution to the three-dimensional obstacle problem converges to the unique solution of a two-dimensional model consisting of a set of variational inequalities that are posed over the abstract completion of a non-empty, closed and convex set. We recall that the two-dimensional limit model obtained departing from Koiter’s model is characterised by the same variational inequalities as the two-dimensional model obtained departing from the variational formulation of a three-dimensional linearly elastic generalised membrane shell subjected to remaining confined in a prescribed half-space, with the remarkable difference that the sets where solutions for these two-dimensional limit models are sought do not coincide, in general. In order to complete the justification of Koiter’s model for linearly elastic generalised membrane shells of the first kind subjected to remaining confined in a half-space, we give sufficient conditions ensuring that the sets where solutions for these two-dimensional limit models are sought coincide.
Keywords. Variational Inequalities $\cdot$ Generalised Membrane Shells $\cdot$ Asymptotic Analysis $\cdot$ Obstacle Problems
MSC 2020. 35J86, 47H07, 74B05.
1. Introduction
The contact of this elastic films with rigid foundations arises in many applicative fields such as Engineering, Material Science, Medicine and Biology. For instance, the description of the motion inside the human heart of the three Aorta valves, which can be regarded as linearly elastic shells, is governed by a mathematical model built up in a way such that each valve remains confined in a certain portion of space without penetrating or being penetrated by the other two valves. In this direction we cite the recent references [46], [47] and [56].
The mechanical deformation of a linearly elastic shell is modelled, in general, via the classical three-dimensional energy for linearised elasticity (cf., e.g., [7]). However, this model has some drawbacks, as it cannot be used to study models where the shell under consideration is made of a non-homogeneous or anisotropic material (cf., e.g., [4] and [5]), or its thickness varies periodically (cf., e.g., [52] and [53]).
To overcome the intrinsic limitations of this model, W.T. Koiter devised, in the papers [29, 30], an alternative model that was meant to provide information on the deformation of a linearly elastic shell by solely considering the deformation of its middle surface. Another important feature of Koiter’s model is that it is valid for all the main types of linearly elastic shells.
The rigorous mathematical justification of Koiter’s model for linearly elastic shells in the case where no obstacles are taken into account was established by P.G. Ciarlet and his associates in the papers [11, 12, 13, 14, 15, 18]. We also mention the papers [36, 37, 38, 42, 49, 50, 51], which are about the numerical analysis and simulations of problems arising in shell theory. For the justification of Koiter’s model in the time-dependent case where the action of temperature is considered and no obstacles are taken into account, we refer to [39].
The study of obstacle problems for thin elastic films was initiated by A. Léger and B. Miara in the papers [32, 33], where the authors focused on linearly elastic shallow shells and linearly elastic flexural shells subjected to a a constraint for the transverse component of the deformation, imposing that the deformed reference configuration must remain confined in a prescribed half-space whose boundary is a plane orthogonal to the vector $\bm{e}^{3}$ of the canonical basis. The choice for the confinement condition adopted by Léger & Miara, however, is not suitable for describing multi-physics problems where a number of shells subjected to a constraint is considered at once.
The papers [16, 17, 19, 20, 40, 41] contribute to making progress in this direction by considering a more general confinement condition, as they discuss the justification of Koiter’s model in the case where the shell, regardless it is an elliptic membrane or a flexural shells, is subjected to remain confined in a general half-space. The confinement condition considered in the aforementioned papers is different, in general, from the usual Signorini condition that is favoured in many textbooks on contact mechanics (cf., e.g., [27]). Indeed, the Signorini condition solely requires the boundary of the deformed three-dimensional elastic body to obey the constraint whereas the confinement condition we will be considering requires that all the points of the deformed reference configuration must obey the constraint. The latter approach seems to be more amenable in the context of a dimension-reduction analysis, where we expect the limit model, defined over a surface in the Euclidean space, to obey the constraint at all the points of one such surface.
To our best knowledge, the justification of Koiter’s model for linearly elastic generalised membrane shells appears to be an unanswered question, to-date. The purpose of this paper is exactly to make progress in this direction. In order to achieve this goal, we will first perform a rigorous asymptotic analysis departing from the classical energy of three-dimensional linearised elasticity for which the admissible competitors to the role of minimiser are sought in a non-empty, closed, and convex subspace of a suitable space, that takes into account the constraint according to which the shell must remain confined in a prescribed half-space.
We will recover the same two-dimensional limit-model as the one recovered in [16], although the set of admissible solutions for the two-dimensional limit-model here recovered is in general larger than the set identified in [16]. We then identify reasonable sufficient conditions for which the two sets coincide. The latter property is established via a density argument.
We will also show in passing that the confinement condition we are considering is actually equivalent to the Signorini condition, provided that the Kirchhoff-Love assumptions are in place.
For the sake of completeness, we also mention the recent papers [44] and [45], that provide examples of variational-inequalities-based models arising in biology, as well as the paper [43], that discusses the modelling and analysis of the melting of a shallow ice sheet by means of a set of doubly non-linear parabolic variational inequalities. We also mention the recent works [23, 31] that explore new ideas in contact problems in non-linear elasticity.
The paper is divided into six sections (including this one). In section 2 we recall some background and notation. In section 3 we formulate a three-dimensional obstacle problem for a general linearly elastic shell. In section 4 we consider - more specifically - the case where the linearly elastic shell under consideration is a linearly elastic generalised membrane of the first kind. Then, we scale the three-dimensional obstacle problem in a way such that the integration domain becomes independent of the thickness parameter. In section 5, a rigorous asymptotic analysis is carried out and the sought set of abstract two-dimensional variational inequalities is recovered. We also observe that the concept of solution for the two-dimensional limit model recovered as a result of the previous asymptotic analysis is different from the concept of solution for the two-dimensional limit model recovered upon completion of a rigorous asymptotic analysis departing from Koiter’s model subjected to the same confinement condition as the one considered in this paper, even though the abstract variational inequalities are the same. In section 6, we give sufficient conditions ensuring that the concept of solution for the two-dimensional limit model recovered departing from the three-dimensional obstacle problem for linearly elastic generalised membrane shells of the first kind and the concept of solution for the same two-dimensional limit model recovered departing from Koiter’s model subjected to the same confinement condition as the one considered in this paper are actually identical.
2. Geometrical preliminaries
For details about the classical notions of differential geometry used in this section and the next one, see, e.g. [8] or [9].
Greek indices, except $\varepsilon$, take their values in the set $\{1,2\}$, while Latin indices, except when they are used for indexing sequences, take their values in the set $\{1,2,3\}$, and the summation convention with respect to repeated indices is systematically used in conjunction with these two rules. The notation $\mathbb{E}^{3}$ designates the three-dimensional Euclidean space whose origin is denoted by $O$. The Euclidean inner product and the vector product of $\bm{u},\bm{v}\in\mathbb{E}^{3}$ are denoted $\bm{u}\cdot\bm{v}$ and $\bm{u}\wedge\bm{v}$, respectively. The Euclidean norm of $\bm{u}\in\mathbb{E}^{3}$ is denoted $|\bm{u}|$. The notation $\delta^{j}_{i}$ designates the Kronecker symbol.
Given an open subset $\Omega$ of $\mathbb{R}^{n}$, notations such as $L^{2}(\Omega)$, $H^{1}(\Omega)$, or $H^{2}(\Omega)$, denote the standard Lebesgue and Sobolev spaces, and the notation $\mathcal{D}(\Omega)$ denotes the space of all functions that are infinitely differentiable over $\Omega$ and have compact supports in $\Omega$. The notation $\|\cdot\|_{X}$ designates the norm of a normed vector space $X$. Given a vector space $X$, the corresponding product space $X\times X\times X$ is denoted by $\bm{X}$.
The abbreviations “a.a.” and “a.e.” mean almost all and almost everywhere, respectively.
The boundary $\Gamma$ of an open subset $\Omega$ in $\mathbb{R}^{n}$ is said to be Lipschitz-continuous if the following conditions are satisfied (cf., e.g., Section 1.18 of [10]): Given an integer $s\geq 1$, there exist constants $\alpha_{1}>0$ and $L>0$, and a finite number of local coordinate systems, with coordinates $\bm{\phi}^{\prime}_{r}=(\phi_{1}^{r},\dots,\phi_{n-1}^{r})\in\mathbb{R}^{n-1}$ and $\phi_{r}=\phi_{n}^{r}$, sets $\tilde{\omega}_{r}:=\{\bm{\phi}_{r}\in\mathbb{R}^{n-1};|\bm{\phi}_{r}|<\alpha_{1}\}$, $1\leq r\leq s$, and corresponding functions
$\tilde{\theta}_{r}:\tilde{\omega}_{r}\to\mathbb{R},$
such that
$\Gamma=\bigcup_{r=1}^{s}\{(\bm{\phi}^{\prime}_{r},\phi_{r});\bm{\phi}^{\prime}_{r}\in\tilde{\omega}_{r}\textup{ and }\phi_{r}=\tilde{\theta}_{r}(\bm{\phi}^{\prime}_{r})\},$
and
$|\tilde{\theta}_{r}(\bm{\phi}^{\prime}_{r})-\tilde{\theta}_{r}(\bm{\upsilon}^{\prime}_{r})|\leq L|\bm{\phi}^{\prime}_{r}-\bm{\upsilon}^{\prime}_{r}|,\textup{ for all }\bm{\phi}^{\prime}_{r},\bm{\upsilon}^{\prime}_{r}\in\tilde{\omega}_{r},\textup{ and all }1\leq r\leq s.$
We observe that the second last formula takes into account overlapping local charts, while the last set of inequalities express the Lipschitz continuity of the mappings $\tilde{\theta}_{r}$.
An open set $\Omega$ is said to be locally on the same side of its boundary $\Gamma$ if, in addition, there exists a constant $\alpha_{2}>0$ such that
$\{(\bm{\phi}^{\prime}_{r},\phi_{r});\bm{\phi}^{\prime}_{r}\in\tilde{\omega}_{r}\textup{ and }\tilde{\theta}_{r}(\bm{\phi}^{\prime}_{r})<\phi_{r}<\tilde{\theta}_{r}(\bm{\phi}^{\prime}_{r})+\alpha_{2}\}$ $\subset\Omega,\quad\textup{ for all }1\leq r\leq s,$
$\{(\bm{\phi}^{\prime}_{r},\phi_{r});\bm{\phi}^{\prime}_{r}\in\tilde{\omega}_{r}\textup{ and }\tilde{\theta}_{r}(\bm{\phi}^{\prime}_{r})-\alpha_{2}<\phi_{r}<\tilde{\theta}_{r}(\bm{\phi}^{\prime}_{r})\}$ $\subset\mathbb{R}^{n}\setminus\overline{\Omega},\quad\textup{ forall }1\leq r\leq s.$
Following [10], a domain in $\mathbb{R}^{n}$ is a bounded Lipschitz domain, namely, a bounded and connected open subset $\Omega$ of $\mathbb{R}^{n}$, whose boundary $\partial\Omega$ is Lipschitz-continuous, the set $\Omega$ being locally on a single side of $\partial\Omega$. We denote by $\partial_{\nu}$ the normal derivative along the boundary $\partial\Omega$.
Let $\omega$ be a domain in $\mathbb{R}^{2}$, let $y=(y_{\alpha})$ denote a generic point in $\overline{\omega}$, and let $\partial_{\alpha}:=\partial/\partial y_{\alpha}$ and $\partial_{\alpha\beta}:=\partial^{2}/\partial y_{\alpha}\partial y_{\beta}$. A mapping $\bm{\theta}\in\mathcal{C}^{1}(\overline{\omega};\mathbb{E}^{3})$ is an immersion if the two vectors
$\bm{a}_{\alpha}(y):=\partial_{\alpha}\bm{\theta}(y)$
are linearly independent at each point $y\in\overline{\omega}$. Then the image $\bm{\theta}(\overline{\omega})$ of the set $\overline{\omega}$ under the mapping $\bm{\theta}$ is a surface in $\mathbb{E}^{3}$, equipped with $y_{1},y_{2}$ as its curvilinear coordinates. Given any point $y\in\overline{\omega}$, the vectors $\bm{a}_{\alpha}(y)$ span the tangent plane to the surface $\bm{\theta}(\overline{\omega})$ at the point $\bm{\theta}(y)$, the unit vector
$\bm{a}_{3}(y):=\frac{\bm{a}_{1}(y)\wedge\bm{a}_{2}(y)}{|\bm{a}_{1}(y)\wedge\bm{a}_{2}(y)|}$
is normal to $\bm{\theta}(\overline{\omega})$ at $\bm{\theta}(y)$, the three vectors $\bm{a}_{i}(y)$ form the covariant basis at $\bm{\theta}(y)$, and the three vectors $\bm{a}^{j}(y)$ defined by the relations
$\bm{a}^{j}(y)\cdot\bm{a}_{i}(y)=\delta^{j}_{i}$
form the contravariant basis at $\bm{\theta}(y)$; note that the vectors $\bm{a}^{\beta}(y)$ also span the tangent plane to $\bm{\theta}(\overline{\omega})$ at $\bm{\theta}(y)$ and that $\bm{a}^{3}(y)=\bm{a}_{3}(y)$.
The first fundamental form of the surface $\bm{\theta}(\overline{\omega})$ is then defined by means of its covariant components
$a_{\alpha\beta}:=\bm{a}_{\alpha}\cdot\bm{a}_{\beta}=a_{\beta\alpha}\in\mathcal{C}^{0}(\overline{\omega}),$
or by means of its contravariant components
$a^{\alpha\beta}:=\bm{a}^{\alpha}\cdot\bm{a}^{\beta}=a^{\beta\alpha}\in\mathcal{C}^{0}(\overline{\omega}).$
Note that the symmetric matrix field $(a^{\alpha\beta})$ is then the inverse of the matrix field $(a_{\alpha\beta})$, that $\bm{a}^{\beta}=a^{\alpha\beta}\bm{a}_{\alpha}$ and $\bm{a}_{\alpha}=a_{\alpha\beta}\bm{a}^{\beta}$, and that the area element along $\bm{\theta}(\overline{\omega})$ is given at each point $\bm{\theta}(y)$, for each $y\in\overline{\omega}$, by $\sqrt{a(y)}\,\mathrm{d}y$, where
$a:=\det(a_{\alpha\beta})\in\mathcal{C}^{0}(\overline{\omega}).$
Given an immersion $\bm{\theta}\in\mathcal{C}^{2}(\overline{\omega};\mathbb{E}^{3})$, the second fundamental form of the surface $\bm{\theta}(\overline{\omega})$ is defined by means of its covariant components
$b_{\alpha\beta}:=\partial_{\alpha}\bm{a}_{\beta}\cdot\bm{a}_{3}=-\bm{a}_{\beta}\cdot\partial_{\alpha}\bm{a}_{3}=b_{\beta\alpha}\in\mathcal{C}^{0}(\overline{\omega}),$
or by means of its mixed components
$b^{\beta}_{\alpha}:=a^{\beta\sigma}b_{\alpha\sigma}\in\mathcal{C}^{0}(\overline{\omega}),$
and the Christoffel symbols associated with the immersion $\bm{\theta}$ are defined by
$\Gamma^{\sigma}_{\alpha\beta}:=\partial_{\alpha}\bm{a}_{\beta}\cdot\bm{a}^{\sigma}=\Gamma^{\sigma}_{\beta\alpha}\in\mathcal{C}^{0}(\overline{\omega}).$
Given an injective immersion $\bm{\theta}\in\mathcal{C}^{2}(\overline{\omega};\mathbb{E}^{3})$ and a vector field $\bm{\eta}=(\eta_{i})\in\mathcal{C}^{1}(\overline{\omega};\mathbb{R}^{3})$, the vector field
$\tilde{\bm{\eta}}:=\eta_{i}\bm{a}^{i}$
can be viewed as a displacement field of the surface $\bm{\theta}(\overline{\omega})$, thus defined by means of its covariant components $\eta_{i}$ over the vectors $\bm{a}^{i}$ of the contravariant bases along the surface. If the norms $|\eta_{i}|_{\mathcal{C}^{1}(\overline{\omega})}$ are small enough, the mapping $(\bm{\theta}+\eta_{i}\bm{a}^{i})\in\mathcal{C}^{1}(\overline{\omega};\mathbb{E}^{3})$ is also an injective immersion, so that the set $(\bm{\theta}+\eta_{i}\bm{a}^{i})(\overline{\omega})$ is also a surface in $\mathbb{E}^{3}$, equipped with the same curvilinear coordinates as those of the surface $\bm{\theta}(\overline{\omega})$, called the deformed surface corresponding to the displacement field $\tilde{\bm{\eta}}=\eta_{i}\bm{a}^{i}$. One can then define the first fundamental form of the deformed surface by means of its covariant components
$a_{\alpha\beta}(\bm{\eta}):=(\bm{a}_{\alpha}+\partial_{\alpha}\tilde{\bm{\eta}})\cdot(\bm{a}_{\beta}+\partial_{\beta}\tilde{\bm{\eta}})a_{\alpha\beta}+\bm{a}_{\alpha}\cdot\partial_{\beta}\tilde{\bm{\eta}}+\partial_{\alpha}\tilde{\bm{\eta}}\cdot\bm{a}_{\beta}+\partial_{\alpha}\tilde{\bm{\eta}}\cdot\partial_{\beta}\tilde{\bm{\eta}}.$
The linear part with respect to $\tilde{\bm{\eta}}$ in the difference $\frac{1}{2}(a_{\alpha\beta}(\bm{\eta})-a_{\alpha\beta})$ is called the linearised change of metric tensor associated with the displacement field $\eta_{i}\bm{a}^{i}$, the covariant components of which are thus defined by
$\gamma_{\alpha\beta}(\bm{\eta}):=\dfrac{1}{2}\left(\bm{a}_{\alpha}\cdot\partial_{\beta}\tilde{\bm{\eta}}+\partial_{\alpha}\tilde{\bm{\eta}}\cdot\bm{a}_{\beta}\right)=\dfrac{1}{2}(\partial_{\beta}\eta_{\alpha}+\partial_{\alpha}\eta_{\beta})-\Gamma^{\sigma}_{\alpha\beta}\eta_{\sigma}-b_{\alpha\beta}\eta_{3}=\gamma_{\beta\alpha}(\bm{\eta}).$
The linear part with respect to $\tilde{\bm{\eta}}$ in the difference $(b_{\alpha\beta}(\bm{\eta})-b_{\alpha\beta})$ is called the linearised change of curvature tensor associated with the displacement field $\eta_{i}\bm{a}^{i}$, the covariant components of which are thus defined by
$\rho_{\alpha\beta}(\bm{\eta})$ $:=(\partial_{\alpha\beta}\tilde{\bm{\eta}}-\Gamma_{\alpha\beta}^{\sigma}\partial_{\sigma}\tilde{\bm{\eta}})\cdot\bm{a}_{3}=\rho_{\beta\alpha}(\bm{\eta})$
$=\partial_{\alpha\beta}\eta_{3}-\Gamma_{\alpha\beta}^{\sigma}\partial_{\sigma}\eta_{3}-b_{\alpha}^{\sigma}b_{\sigma\beta}\eta_{3}+b_{\alpha}^{\sigma}(\partial_{\beta}\eta_{\sigma}-\Gamma_{\beta\sigma}^{\tau}\eta_{\tau})+b_{\beta}^{\tau}(\partial_{\alpha}\eta_{\tau}-\Gamma_{\alpha\tau}^{\sigma}\eta_{\sigma})+(\partial_{\alpha}b_{\beta}^{\tau}+\Gamma_{\alpha\sigma}^{\tau}b_{\beta}^{\sigma}-\Gamma_{\alpha\beta}^{\sigma}b_{\sigma}^{\tau})\eta_{\tau}.$
3. A general three-dimensional obstacle problem for a linearly elastic shell
Let $\omega$ be a domain in $\mathbb{R}^{2}$, let $\gamma:=\partial\omega$, and let $\gamma_{0}$ be a non-empty relatively open subset of $\gamma$. For each $\varepsilon>0$, we define the sets
$\Omega^{\varepsilon}:=\omega\times(-\varepsilon,\varepsilon)\quad\textup{ and }\quad\Gamma^{\varepsilon}_{0}:=\gamma_{0}\times[-\varepsilon,\varepsilon],$
we let $x^{\varepsilon}=(x^{\varepsilon}_{i})$ designate a generic point in the set $\overline{\Omega^{\varepsilon}}$, and we let $\partial^{\varepsilon}_{i}:=\partial/\partial x^{\varepsilon}_{i}$. Hence we also have $x^{\varepsilon}_{\alpha}=y_{\alpha}$ and $\partial^{\varepsilon}_{\alpha}=\partial_{\alpha}$.
Given an injective immersion $\bm{\theta}\in\mathcal{C}^{3}(\overline{\omega};\mathbb{E}^{3})$ and $\varepsilon>0$, consider a shell with middle surface $\bm{\theta}(\overline{\omega})$ and with constant thickness $2\varepsilon$. This means that the reference configuration of the shell is the set $\bm{\Theta}^{\varepsilon}(\overline{\Omega^{\varepsilon}})$, where the mapping $\bm{\Theta}^{\varepsilon}:\overline{\Omega^{\varepsilon}}\to\mathbb{E}^{3}$ is defined by
$\bm{\Theta}^{\varepsilon}(x^{\varepsilon}):=\bm{\theta}(y)+x^{\varepsilon}_{3}\bm{a}^{3}(y),\quad\textup{ for all }x^{\varepsilon}=(y,x^{\varepsilon}_{3})\in\overline{\Omega^{\varepsilon}}.$
One can then show (cf., e.g., Theorem 3.1-1 of [8]) that, if $\varepsilon>0$ is small enough, such a mapping $\bm{\Theta}^{\varepsilon}\in\mathcal{C}^{2}(\overline{\Omega^{\varepsilon}};\mathbb{E}^{3})$ is also an injective immersion, in the sense that the three vectors
$\bm{g}^{\varepsilon}_{i}(x^{\varepsilon}):=\partial^{\varepsilon}_{i}\bm{\Theta}^{\varepsilon}(x^{\varepsilon}),$
are linearly independent at each point $x^{\varepsilon}\in\overline{\Omega^{\varepsilon}}$. The vectors $\{\bm{g}^{\varepsilon}_{i}\}_{i=1}^{3}$ then constitute the covariant basis at the point $\bm{\Theta}^{\varepsilon}(x^{\varepsilon})$, while the three vectors $\bm{g}^{j,\varepsilon}(x^{\varepsilon})$ defined by the relations
$\bm{g}^{j,\varepsilon}(x^{\varepsilon})\cdot\bm{g}^{\varepsilon}_{i}(x^{\varepsilon})=\delta^{j}_{i},$
constitute the contravariant basis at the same point. It will be implicitly assumed in the sequel that $\varepsilon>0$ is small enough so that $\bm{\Theta}^{\varepsilon}:\overline{\Omega^{\varepsilon}}\to\mathbb{E}^{3}$ is an injective immersion.
One then defines the metric tensor associated with the immersion $\bm{\Theta}^{\varepsilon}$ by means of its covariant components
$g^{\varepsilon}_{ij}:=\bm{g}^{\varepsilon}_{i}\cdot\bm{g}^{\varepsilon}_{j}\in\mathcal{C}^{1}(\overline{\Omega^{\varepsilon}}),$
or by means of its contravariant components
$g^{ij,\varepsilon}:=\bm{g}^{i,\varepsilon}\cdot\bm{g}^{j,\varepsilon}\in\mathcal{C}^{1}(\overline{\Omega^{\varepsilon}}).$
Note that the symmetric matrix field $(g^{ij,\varepsilon})$ is then the inverse of the matrix field $(g^{\varepsilon}_{ij})$, that $\bm{g}^{j,\varepsilon}=g^{ij,\varepsilon}\bm{g}^{\varepsilon}_{i}$ and $\bm{g}^{\varepsilon}_{i}=g^{\varepsilon}_{ij}\bm{g}^{j,\varepsilon}$, and that the volume element in $\bm{\Theta}^{\varepsilon}(\overline{\Omega^{\varepsilon}})$ is given at each point $\bm{\Theta}^{\varepsilon}(x^{\varepsilon})$, for all $x^{\varepsilon}\in\overline{\Omega^{\varepsilon}}$, by $\sqrt{g^{\varepsilon}(x^{\varepsilon})}\,\mathrm{d}x^{\varepsilon}$, where
$g^{\varepsilon}:=\det(g^{\varepsilon}_{ij})\in\mathcal{C}^{1}(\overline{\Omega^{\varepsilon}}).$
One also defines the Christoffel symbols associated with the immersion $\bm{\Theta}^{\varepsilon}$ by
$\Gamma^{p,\varepsilon}_{ij}:=\partial_{i}^{\varepsilon}\bm{g}^{\varepsilon}_{j}\cdot\bm{g}^{p,\varepsilon}=\Gamma^{p,\varepsilon}_{ji}\in\mathcal{C}^{0}(\overline{\Omega^{\varepsilon}}).$
Note that $\Gamma^{3,\varepsilon}_{\alpha 3}=\Gamma^{p,\varepsilon}_{33}=0$.
Given a vector field $\bm{v}^{\varepsilon}=(v^{\varepsilon}_{i})\in\mathcal{C}^{1}(\overline{\Omega^{\varepsilon}};\mathbb{R}^{3})$, the associated vector field
$\tilde{\bm{v}}^{\varepsilon}:=v^{\varepsilon}_{i}\bm{g}^{i,\varepsilon},$
can be viewed as a displacement field of the reference configuration $\bm{\Theta}^{\varepsilon}(\overline{\Omega^{\varepsilon}})$ of the shell, thus defined by means of its covariant components $v^{\varepsilon}_{i}$ over the vectors $\bm{g}^{i,\varepsilon}$ of the contravariant bases in the reference configuration.
If the norms $\|v^{\varepsilon}_{i}\|_{\mathcal{C}^{1}(\overline{\Omega^{\varepsilon}})}$ are small enough, the mapping $(\bm{\Theta}^{\varepsilon}+v^{\varepsilon}_{i}\bm{g}^{i,\varepsilon})$ is also an injective immersion, so that one can also define the metric tensor of the deformed configuration $(\bm{\Theta}^{\varepsilon}+v^{\varepsilon}_{i}\bm{g}^{i,\varepsilon})(\overline{\Omega^{\varepsilon}})$ by means of its covariant components
$g^{\varepsilon}_{ij}(v^{\varepsilon}):=(\bm{g}^{\varepsilon}_{i}+\partial^{\varepsilon}_{i}\tilde{\bm{v}}^{\varepsilon})\cdot(\bm{g}^{\varepsilon}_{j}+\partial^{\varepsilon}_{j}\tilde{\bm{v}}^{\varepsilon})=g^{\varepsilon}_{ij}+\bm{g}^{\varepsilon}_{i}\cdot\partial_{j}^{\varepsilon}\tilde{\bm{v}}^{\varepsilon}+\partial^{\varepsilon}_{i}\tilde{\bm{v}}^{\varepsilon}\cdot\bm{g}^{\varepsilon}_{j}+\partial_{i}^{\varepsilon}\tilde{\bm{v}}^{\varepsilon}\cdot\partial_{j}^{\varepsilon}\tilde{\bm{v}}^{\varepsilon}.$
The linear part with respect to $\tilde{\bm{v}}^{\varepsilon}$ in the difference $\dfrac{1}{2}(g^{\varepsilon}_{ij}(\bm{v}^{\varepsilon})-g^{\varepsilon}_{ij})$ is then called the linearised strain tensor associated with the displacement field $v^{\varepsilon}_{i}\bm{g}^{i,\varepsilon}$, the covariant components of which are thus defined by
$e^{\varepsilon}_{i\|j}(\bm{v}^{\varepsilon}):=\frac{1}{2}\left(\bm{g}^{\varepsilon}_{i}\cdot\partial_{j}^{\varepsilon}\tilde{\bm{v}}^{\varepsilon}+\partial^{\varepsilon}_{i}\tilde{\bm{v}}^{\varepsilon}\cdot\bm{g}^{\varepsilon}_{j}\right)=\frac{1}{2}(\partial^{\varepsilon}_{j}v^{\varepsilon}_{i}+\partial^{\varepsilon}_{i}v^{\varepsilon}_{j})-\Gamma^{p,\varepsilon}_{ij}v^{\varepsilon}_{p}=e_{j\|i}^{\varepsilon}(\bm{v}^{\varepsilon}).$
The functions $e^{\varepsilon}_{i\|j}(\bm{v}^{\varepsilon})$ are called the linearised strains in curvilinear coordinates associated with the displacement field $v^{\varepsilon}_{i}\bm{g}^{i,\varepsilon}$.
We assume throughout this paper that, for each $\varepsilon>0$, the reference configuration $\bm{\Theta}^{\varepsilon}(\overline{\Omega^{\varepsilon}})$ of the shell is a natural state (i.e., stress-free) and that the material constituting the shell is homogeneous, isotropic, and linearly elastic. The behaviour of such an elastic material is thus entirely governed by its two Lamé constants $\lambda\geq 0$ and $\mu>0$ (for details, see, e.g., Section 3.8 of [7]).
We will also assume that the shell is subjected to applied body forces whose density per unit volume is defined by means of its covariant components $f^{i,\varepsilon}\in L^{2}(\Omega^{\varepsilon})$, and to a homogeneous boundary condition of place along the portion $\Gamma^{\varepsilon}_{0}$ of its lateral face (i.e., the displacement vanishes on $\Gamma^{\varepsilon}_{0}$).
The confinement condition considered in this paper, which was originally suggested by Brezis & Stampacchia [3], does not take any traction forces into account. One of the reasons for the latter choice is that obstacle problems are free-boundary problems and, by Classical Mechanics, there could be no traction forces acting on the portion of the boundary of the three-dimensional shell that engages contact with the obstacle.
In this paper we consider a specific obstacle problem for such a shell, in the sense that the shell is subjected to a confinement condition, expressing that any admissible displacement vector field $v^{\varepsilon}_{i}\bm{g}^{i,\varepsilon}$ must be such that all the points of the corresponding deformed configuration remain in a prescribed half-space of the form
:=$\mathbb{H}:=\{x\in\mathbb{E}^{3};\,\bm{Ox}\cdot\bm{q}\geq 0\},$
where $\bm{q}$ is a unit-vector which is given once and for all. It is evident that the vector $\bm{q}$ is thus orthogonal to the plane associated with the half-space where the linearly elastic shell is required to remain confined. In other words, any admissible displacement field must satisfy
$\left(\bm{\Theta}^{\varepsilon}(x^{\varepsilon})+v^{\varepsilon}_{i}(x^{\varepsilon})\bm{g}^{i,\varepsilon}(x^{\varepsilon})\right)\cdot\bm{q}\geq 0,$
for all $x^{\varepsilon}\in\overline{\Omega^{\varepsilon}}$, or possibly only for almost all $x^{\varepsilon}\in\Omega^{\varepsilon}$ when the covariant components $v^{\varepsilon}_{i}$ are required to belong to the Sobolev space $H^{1}(\Omega^{\varepsilon})$ as in Theorem 3.1 below.
We will of course assume that the reference configuration satisfies the confinement condition, i.e., that
$\bm{\Theta}^{\varepsilon}(\overline{\Omega^{\varepsilon}})\subset\mathbb{H}.$
It is to be emphasised that the above confinement condition considerably departs from the usual Signorini condition favored by most authors, who usually require that only the points of the undeformed and deformed “lower face” $\omega\times\{-\varepsilon\}$ of the reference configuration satisfy the confinement condition (see, e.g., [32], [33], [48]). Clearly, the confinement condition considered in the present paper, which is inspired by the formulation proposed by Brézis & Stampacchia [3], is more physically realistic, since a Signorini condition imposed only on the lower face of the reference configuration does not prevent – at least “mathematically” – other points of the deformed reference configuration to “cross” the plane $\{x\in\mathbb{E}^{3};\;\bm{Ox}\cdot\bm{q}=0\}$ and then to end up on the “other side” of this plane (cf., e.g., Chapter 63 in [55]).
Such a confinement condition renders the asymptotic analysis considerably more difficult, however, as the constraint now bears on a vector field, the displacement vector field of the reference configuration, instead of on only a single component of this field.
The mathematical modelling of such an obstacle problem for a linearly elastic shell thus follow straightforwardly. Indeed, apart from the confinement condition, the rest, i.e., the function space and the expression of the quadratic energy $J^{\varepsilon}$, is classical (see, e.g. [8]). More specifically, let
$A^{ijk\ell,\varepsilon}:=\lambda g^{ij,\varepsilon}g^{k\ell,\varepsilon}+\mu\left(g^{ik,\varepsilon}g^{j\ell,\varepsilon}+g^{i\ell,\varepsilon}g^{jk,\varepsilon}\right)=A^{jik\ell,\varepsilon}=A^{k\ell ij,\varepsilon},$
denote the contravariant components of the elasticity tensor of the elastic material constituting the shell. Then the unknown of the problem, which is the vector field $\bm{u}^{\varepsilon}=(u^{\varepsilon}_{i})$ where the functions $u^{\varepsilon}_{i}:\overline{\Omega^{\varepsilon}}\to\mathbb{R}$ are the three covariant components of the unknown “three-dimensional” displacement vector field $u^{\varepsilon}_{i}\bm{g}^{i,\varepsilon}$ of the reference configuration of the shell, should minimise the energy $J^{\varepsilon}:\bm{H}^{1}(\Omega^{\varepsilon})\to\mathbb{R}$ defined by
$J^{\varepsilon}(\bm{v}^{\varepsilon}):=\frac{1}{2}\int_{\Omega^{\varepsilon}}A^{ijk\ell,\varepsilon}e^{\varepsilon}_{k\|\ell}(\bm{v}^{\varepsilon})e^{\varepsilon}_{i\|j}(\bm{v}^{\varepsilon})\sqrt{g^{\varepsilon}}\,\mathrm{d}x^{\varepsilon}-\int_{\Omega^{\varepsilon}}f^{i,\varepsilon}v^{\varepsilon}_{i}\sqrt{g^{\varepsilon}}\,\mathrm{d}x^{\varepsilon},$
for each $\bm{v}^{\varepsilon}=(v^{\varepsilon}_{i})\in\bm{H}^{1}(\Omega^{\varepsilon})$ over the set of admissible displacements defined by:
$\bm{U}(\Omega^{\varepsilon}):=\{\bm{v}^{\varepsilon}=(v^{\varepsilon}_{i})\in\bm{H}^{1}(\Omega^{\varepsilon});\bm{v}^{\varepsilon}=\textbf{0}\text{ on }\Gamma^{\varepsilon}_{0}\textup{ and }(\bm{\Theta}^{\varepsilon}(x^{\varepsilon})+v^{\varepsilon}_{i}(x^{\varepsilon})\bm{g}^{i,\varepsilon}(x^{\varepsilon}))\cdot\bm{q}\geq 0\textup{ for a.a. }x^{\varepsilon}\in\Omega^{\varepsilon}\}.$
The solution to this minimisation problem exists and is unique, and it can be also characterised as the unique solution of the following problem:
Problem $\mathcal{P}(\Omega^{\varepsilon})$.
Find $\bm{u}^{\varepsilon}=(u_{i}^{\varepsilon})\in\bm{U}(\Omega^{\varepsilon})$ that satisfies the following variational inequalities:
$\int_{\Omega^{\varepsilon}}A^{ijk\ell,\varepsilon}e^{\varepsilon}_{k\|\ell}(\bm{u}^{\varepsilon})\left(e^{\varepsilon}_{i\|j}(\bm{v}^{\varepsilon})-e^{\varepsilon}_{i\|j}(\bm{u}^{\varepsilon})\right)\sqrt{g^{\varepsilon}}\,\mathrm{d}x^{\varepsilon}\geq\int_{\Omega^{\varepsilon}}f^{i,\varepsilon}(v^{\varepsilon}_{i}-u^{\varepsilon}_{i})\sqrt{g^{\varepsilon}}\,\mathrm{d}x^{\varepsilon},$
for all $\bm{v}^{\varepsilon}=(v^{\varepsilon}_{i})\in\bm{U}(\Omega^{\varepsilon})$. $\blacksquare$
The following result can be thus straightforwardly proved.
Theorem 3.1.
The quadratic minimisation problem: Find a vector field $\bm{u}^{\varepsilon}\in\bm{U}(\Omega^{\varepsilon})$ such that
$J^{\varepsilon}(\bm{u}^{\varepsilon})=\inf_{\bm{v}^{\varepsilon}\in\bm{U}(\Omega^{\varepsilon})}J^{\varepsilon}(\bm{v}^{\varepsilon}),$
has one and only one solution. Besides, $\bm{u}^{\varepsilon}$ is also the unique solution of Problem $\mathcal{P}(\Omega^{\varepsilon})$. ∎
Since $\bm{\theta}(\overline{\omega})\subset\bm{\Theta}^{\varepsilon}(\overline{\Omega^{\varepsilon}})$, it evidently follows that $\bm{\theta}(y)\cdot\bm{q}\geq 0$ for all $y\in\overline{\omega}$. But in fact, a stronger property holds (cf. Lemma 2.1 of [20]).
Lemma 3.1.
Let $\omega$ be a domain in $\mathbb{R}^{2}$, let $\bm{\theta}\in\mathcal{C}^{1}(\overline{\omega};\mathbb{E}^{3})$ be an injective immersion, let $\bm{q}\in\mathbb{E}^{3}$ be a given unit-vector, and let $\varepsilon>0$. Then the inclusion
$\bm{\Theta}^{\varepsilon}(\overline{\Omega^{\varepsilon}})\subset\mathbb{H}=\{x\in\mathbb{E}^{3};\;\bm{Ox}\cdot\bm{q}\geq 0\}$
implies that:
$d:=\min_{y\in\overline{\omega}}(\bm{\theta}(y)\cdot\bm{q})>0.$
4. The scaled three-dimensional problem for a family of generalised membrane shells
In section 3, we considered an obstacle problem for “general” linearly elastic shells. From now on, we will restrict ourselves to a specific class of shells, according to the following definition (cf., e.g., Section 5.1 of [8]): A linearly elastic shell is said to be a linearly elastic generalised membrane shell if the following two additional assumptions are simultaneously satisfied: First, $\text{length}\,\gamma_{0}>0$ (an assumption that is satisfied if $\gamma_{0}$ is a non-empty relatively open subset of $\gamma$, as assumed here). Second, the space of admissible linearised inextensional displacements defined by
$\bm{V}_{F}(\omega):=\{\bm{\eta}=(\eta_{i})\in H^{1}(\omega)\times H^{1}(\omega)\times H^{2}(\omega);\eta_{i}=\partial_{\nu}\eta_{3}=0\textup{ on }\gamma_{0}\textup{ and }\gamma_{\alpha\beta}(\bm{\eta})=0\textup{ in }\omega\},$
does not contain any non-zero vectors, i.e.,
$\bm{V}_{F}(\omega)=\{\bm{0}\},$
but, the shell is not a linearly elastic elliptic membrane shell (note that, in light of the results in [8], the space $\bm{V}_{F}(\omega)$ indeed reduces to $\{\bm{0}\}$ if the shell is a linearly elastic elliptic membrane one). The second condition in the definition of a generalised membrane shell, viz.,
$\bm{V}_{F}(\omega)=\{\bm{0}\},$
is equivalent to stating that the semi-norm $\left|\cdot\right|_{\omega}^{M}$ defined by
$|\bm{\eta}|_{\omega}^{M}:=\left\{\sum_{\alpha,\beta}|\gamma_{\alpha\beta}(\bm{\eta})|_{L^{2}(\omega)}^{2}\right\}^{1/2},$
for each $\bm{\eta}=(\eta_{i})\in H^{1}(\omega)\times H^{1}(\omega)\times L^{2}(\omega)$ becomes a norm over the space
$\bm{V}_{K}(\omega)=\{\bm{\eta}=(\eta_{i})\in H^{1}(\omega)\times H^{1}(\omega)\times H^{2}(\omega);\eta_{i}=\partial_{\nu}\eta_{3}=0\textup{ on }\gamma_{0}\}.$
Generalised membrane shells are themselves subdivided into two categories described in terms of the spaces:
$\bm{V}(\omega)$ :=$:=\{\bm{\eta}=(\eta_{i})\in\bm{H}^{1}(\omega);\bm{\eta}={\bm{0}}\textup{ on }\gamma_{0}\},$
$\bm{V}_{0}(\omega)$ :=$:=\{\bm{\eta}=(\eta_{i})\in\bm{V}(\omega);\gamma_{\alpha\beta}(\bm{\eta})=0\textup{ in }\omega\}.$
A generalised membrane shell is “of the first kind” if
$\bm{V}_{0}(\omega)=\{\bm{0}\},$
or, equivalently, if the semi-norm $\left|\cdot\right|_{\omega}^{M}$ is already a norm over the space $\bm{V}(\omega)$ (hence, a fortiori, over the space $\bm{V}_{K}(\omega)\subset\bm{V}(\omega)$). Otherwise, i.e., if
${\bm{V}_{F}(\omega)=\{\bm{0}\}}\quad\textup{but}\quad{\bm{V}_{0}(\omega)\neq\{\bm{0}\}},$
or, equivalently, if the semi-norm $\left|\cdot\right|_{\omega}^{M}$ is a norm over $\bm{V}_{K}(\omega)$ but not over $\bm{V}(\omega)$, the linearly elastic shell is a generalised membrane shell “of the second kind”. In this paper we shall only consider generalised membranes “of the first kind” as they are the most frequently encountered in the practice.
In this paper, we consider the obstacle problem as defined in section 3 for a family of linearly elastic generalised membrane shells of the first kind, all sharing the same middle surface and whose thickness $02\varepsilon>$ 0 is considered as a “small” parameter approaching zero. Our objective then consists in performing an asymptotic analysis as $\varepsilon\to 0$, so as to seek whether we can identify a two-dimensional limit model. To this end, we shall resort to a (by now standard) methodology first proposed by Ciarlet & Lods in [15] (see also Theorem 5.6-1 of [8]). To begin with, we “scale” Problem $\mathcal{P}(\Omega^{\varepsilon})$, over a stretched domain $\Omega$, using appropriate scalings on the unknowns and assumptions on the data. Define
$\Omega:=\omega\times(-1,1),$
let $x=(x_{i})$ denote a generic point in the set $\overline{\Omega}$, and let $\partial_{i}:=\partial/\partial x_{i}$. With each point $x=(x_{i})\in\overline{\Omega}$, we associate the point $x^{\varepsilon}=(x^{\varepsilon}_{i})$ defined by
$x^{\varepsilon}_{\alpha}:=x_{\alpha}=y_{\alpha}\quad\text{ and }\quad x^{\varepsilon}_{3}:=\varepsilon x_{3},$
so that $\partial^{\varepsilon}_{\alpha}=\partial_{\alpha}$ and $\partial^{\varepsilon}_{3}=\frac{1}{\varepsilon}\partial_{3}$. To the unknown $\bm{u}^{\varepsilon}=(u^{\varepsilon}_{i})$ and to the vector fields $\bm{v}^{\varepsilon}=(v^{\varepsilon}_{i})$ appearing in the formulation of Problem $\mathcal{P}(\Omega^{\varepsilon})$ we then associate the scaled unknown $\bm{u}(\varepsilon)=(u_{i}(\varepsilon))$ and the scaled vector fields $\bm{v}=(v_{i})$ by letting
$u_{i}(\varepsilon)(x):=u^{\varepsilon}_{i}(x^{\varepsilon})\text{ and }v_{i}(x):=v^{\varepsilon}_{i}(x^{\varepsilon}),$
at each $x\in\overline{\Omega}$. Finally, we assume that there exist functions $f^{i}\in L^{2}(\Omega)$ independent on $\varepsilon$ such that the following assumptions on the data hold:
$f^{i,\varepsilon}(x^{\varepsilon})=f^{i}(x),\quad\text{ for a.a. }x\in\Omega.$
Note that the independence on $\varepsilon$ of the Lamé constants assumed in section 3 in the formulation of Problem $\mathcal{P}(\Omega^{\varepsilon})$ implicitly constitutes another assumption on the data.
In view of the proposed scaling, we define the “scaled” version of the geometrical entities introduced in section 2:
$\bm{g}^{i}(\varepsilon)(x)$ $:=\bm{g}^{i,\varepsilon}(x^{\varepsilon}),\quad\textup{ for all }x\in\overline{\Omega},$
$g(\varepsilon)(x)$ $:=g^{\varepsilon}(x^{\varepsilon})\text{ and }A^{ijk\ell}(\varepsilon)(x):=A^{ijk\ell,\varepsilon}(x^{\varepsilon}),\quad\textup{ for all }x\in\overline{\Omega},$
$e_{\alpha\|\beta}(\varepsilon;\bm{v})$ $:=\frac{1}{2}(\partial_{\beta}v_{\alpha}+\partial_{\alpha}v_{\beta})-\Gamma^{k}_{\alpha\beta}(\varepsilon)v_{k}=e_{\beta\|\alpha}(\varepsilon;\bm{v}),$
$e_{\alpha\|3}(\varepsilon;\bm{v})=e_{3\|\alpha}(\varepsilon;\bm{v})$ $:=\frac{1}{2}\left(\frac{1}{\varepsilon}\partial_{3}v_{\alpha}+\partial_{\alpha}v_{3}\right)-\Gamma^{\sigma}_{\alpha 3}(\varepsilon)v_{\sigma},$
$e_{3\|3}(\varepsilon;\bm{v})$ $:=\frac{1}{\varepsilon}\partial_{3}v_{3},$
where
$\Gamma^{p}_{ij}(\varepsilon)(x):=\Gamma^{p,\varepsilon}_{ij}(x^{\varepsilon}),\quad\textup{ for all }x\in\overline{\Omega}.$
Define the space
$\bm{V}(\Omega):=\{\bm{v}=(v_{i})\in\bm{H}^{1}(\Omega);\;\bm{v}=\bm{0}\textup{ on }\gamma_{0}\times(-1,1)\},$
and define, for each $\varepsilon>0$, the set
$\bm{U}(\varepsilon;\Omega)$ :=$:=\{\bm{v}=(v_{i})\in\bm{V}(\Omega);\big{(}\bm{\theta}(y)+\varepsilon x_{3}\bm{a}_{3}(y)+v_{i}(x)\bm{g}^{i}(\varepsilon)(x)\big{)}\cdot\bm{q}\geq 0\textup{ for a.a. }x=(y,x_{3})\in\Omega\}.$
We are thus in a position to introduce the “scaled” version of Problem $\mathcal{P}(\Omega^{\varepsilon})$, that will be denoted in what follows by $\mathcal{P}(\varepsilon;\Omega)$.
Problem $\mathcal{P}(\varepsilon;\Omega)$.
Find $\bm{u}(\varepsilon)=(u_{i}(\varepsilon))\in\bm{U}(\varepsilon;\Omega)$ that satisfies the following variational inequalities:
$\int_{\Omega}A^{ijk\ell}(\varepsilon)e_{k\|\ell}(\varepsilon;\bm{u}(\varepsilon))\left(e_{i\|j}(\varepsilon;\bm{v})-e_{i\|j}(\varepsilon;\bm{u}(\varepsilon))\right)\sqrt{g(\varepsilon)}\,\mathrm{d}x\geq\int_{\Omega}f^{i}(v_{i}-u_{i}(\varepsilon))\sqrt{g(\varepsilon)}\,\mathrm{d}x,$
for all $\bm{v}=(v_{i})\in\bm{U}(\varepsilon;\Omega)$. $\blacksquare$
Theorem 4.1.
The scaled unknown $\bm{u}(\varepsilon)$ is the unique solution of the variational Problem $\mathcal{P}(\varepsilon;\Omega)$.
Proof.
The variational Problem $\mathcal{P}(\varepsilon;\Omega)$ simply constitutes a re-writing of the variational Problem $\mathcal{P}(\Omega^{\varepsilon})$, this time in terms of the scaled unknown $\bm{u}(\varepsilon)$, of the vector fields $\bm{v}$, and of the functions $f^{i}$, which are now all defined over the domain $\Omega$. Then the assertion follows from this observation. ∎
The functions $e_{i\|j}(\varepsilon;\bm{v})$ appearing in Problem $\mathcal{P}(\varepsilon;\Omega)$ are called the scaled linearised strains in curvilinear coordinates associated with the scaled displacement vector field $v_{i}\bm{g}^{i}(\varepsilon)$.
For later purposes (like in Lemma 4.1 below), we also let
$\bm{g}_{i}(\varepsilon)(x):=\bm{g}^{\varepsilon}_{i}(x^{\varepsilon})\textup{ at each }x\in\overline{\Omega}.$
It is immediately verified (cf., e.g., [8]) that other assumptions on the data are possible that would give rise to the same problem over the fixed domain $\Omega$. For instance, should the Lamé constants (now denoted) $\lambda^{\varepsilon}$ and $\mu^{\varepsilon}$ appearing in Problem $\mathcal{P}(\Omega^{\varepsilon})$ be of the form $\lambda^{\varepsilon}=\varepsilon^{t}\lambda$ and $\mu^{\varepsilon}=\varepsilon^{t}\mu$, where $\lambda\geq 0$ and $\mu$ are constants independent of $\varepsilon$ and $t$ is an arbitrary real number, the same Problem $\mathcal{P}(\varepsilon;\Omega)$ arises if we assume that the components of the applied body force density are now of the form
$f^{i,\varepsilon}(x^{\varepsilon})=\varepsilon^{t}f^{i}(x),\text{ for a.a. }x\in\Omega,$
where the functions $f^{i}\in L^{2}(\Omega)$ are independent of $\varepsilon$.
The next lemma assembles various asymptotic properties as $\varepsilon\to 0$ of functions and vector fields appearing in the formulation of Problem $\mathcal{P}(\varepsilon;\Omega)$. These properties will be repeatedly used in the proof of the convergence theorem (Theorem 5.1).
In the statement of the next preparatory lemma (cf., e.g., Theorems 3.3-1 and 3.3-2 of [8]), the notation “$\mathcal{O}(\varepsilon)$”, or “$\mathcal{O}(\varepsilon^{2})$”, stands for a remainder that is of order $\varepsilon$, or $\varepsilon^{2}$, with respect to the sup-norm over the set $\overline{\Omega}$, and any function, or vector-valued function, of the variable $y\in\overline{\omega}$, such as $a^{\alpha\beta},b_{\alpha\beta},\bm{a}^{i}$, etc. (all these are defined in section 2) is identified with the function, or vector-valued function, of $x=(y,x_{3})\in\overline{\Omega}=\overline{\omega}\times[-1,1]$ that takes the same value at $x_{3}=0$ and is independent of $x_{3}\in\left[-1,1\right]$; for brevity, this extension from $\overline{\omega}$ to $\overline{\Omega}$ is designated with the same notation.
Lemma 4.1.
Let $\bm{\theta}\in\mathcal{C}^{3}(\overline{\omega};\mathbb{R}^{3})$ be an injective immersion. Let $\varepsilon_{0}$ be defined as in Theorem 3.1-1 of [8]. The functions $A^{ijk\ell}(\varepsilon)=A^{jik\ell}(\varepsilon)=A^{k\ell ij}(\varepsilon)$ have the following properties:
$A^{ijk\ell}(\varepsilon)=A^{ijk\ell}(0)+\mathcal{O}(\varepsilon),\quad A^{\alpha\beta\sigma 3}(\varepsilon)=A^{\alpha 333}(\varepsilon)=0,$
for all $0<\varepsilon\leq\varepsilon_{0}$, where
$A^{\alpha\beta\sigma\tau}(0)$ $=\lambda a^{\alpha\beta}a^{\sigma\tau}+\mu(a^{\alpha\sigma}a^{\beta\tau}+a^{\alpha\tau}a^{\beta\sigma}),$
$A^{\alpha\beta 33}(0)$ $=\lambda a^{\alpha\beta},\quad A^{\alpha 3\sigma 3}(0)=\mu a^{\alpha\sigma},\quad A^{3333}(0)=\lambda+2\mu,$
and there exists a constant $C_{0}>0$ such that
$\sum_{i,j}|t_{ij}|^{2}\leq C_{0}A^{ijk\ell}(\varepsilon)(x)t_{k\ell}t_{ij},$
for all $0<\varepsilon\leq\varepsilon_{0}$, all $x\in\overline{\Omega}$, and all symmetric matrices $(t_{ij})$.
The functions $\Gamma^{p}_{ij}(\varepsilon)$ and $g(\varepsilon)$ have the following properties:
$\Gamma^{\sigma}_{\alpha\beta}(\varepsilon)$ $=\Gamma^{\sigma}_{\alpha\beta}-\varepsilon x_{3}(\partial_{\alpha}b^{\sigma}_{\beta}+\Gamma^{\sigma}_{\alpha\tau}b^{\tau}_{\beta}-\Gamma^{\tau}_{\alpha\beta}b^{\sigma}_{\tau})+\mathcal{O}(\varepsilon^{2}),$
$\Gamma^{3}_{\alpha\beta}(\varepsilon)$ $=b_{\alpha\beta}-\varepsilon x_{3}b^{\sigma}_{\alpha}b_{\sigma\beta},$
$\partial_{3}\Gamma^{p}_{\alpha\beta}(\varepsilon)$ $=\mathcal{O}(\varepsilon),$
$\Gamma^{\sigma}_{\alpha 3}(\varepsilon)$ $=-b^{\sigma}_{\alpha}-\varepsilon x_{3}b^{\tau}_{\alpha}b^{\sigma}_{\tau}+\mathcal{O}(\varepsilon^{2}),$
$\Gamma^{3}_{\alpha 3}(\varepsilon)$ $=\Gamma^{p}_{33}(\varepsilon)=0,$
$g(\varepsilon)$ $=a+\mathcal{O}(\varepsilon),$
for all $0<\varepsilon\leq\varepsilon_{0}$ and all $x\in\overline{\Omega}$. In particular then, there exist constants $g_{0}$ and $g_{1}$ such that
$0<g_{0}\leq g(\varepsilon)(x)\leq g_{1},\quad\textup{ for all }0<\varepsilon\leq\varepsilon_{0}\textup{ and all }x\in\overline{\Omega}.$
The vector fields $\bm{g}_{i}(\varepsilon)$ and $\bm{g}^{j}(\varepsilon)$ have the following properties:
$\bm{g}_{\alpha}(\varepsilon)$ $=\bm{a}_{\alpha}-\varepsilon x_{3}b^{\sigma}_{\alpha}\bm{a}_{\sigma},$
$\bm{g}^{\alpha}(\varepsilon)$ $=\bm{a}^{\alpha}+\varepsilon x_{3}b^{\alpha}_{\sigma}\bm{a}^{\sigma}+\mathcal{O}(\varepsilon^{2}),$
$\bm{g}_{3}(\varepsilon)$ $=\bm{a}_{3}=\bm{a}^{3}=\bm{g}^{3}(\varepsilon).$
We recall (cf., e.g., [8]) that the various relations and estimates in Lemma 4.1 hold in fact for any family of linearly elastic shell.
The next result, which is of independent interest and is not exploited in the forthcoming asymptotic analyses, relates the confinement condition we are considering in this paper with the classical Signorini condition (cf., e.g., Chapter 6 in [27]), according to which:
$(\bm{\Theta}(y,x_{3}^{\varepsilon})+v_{i}^{\varepsilon}(y,x_{3}^{\varepsilon})\bm{g}^{i,\varepsilon}(y,x_{3}^{\varepsilon}))\cdot\bm{q}\geq 0,\quad\textup{ for a.a. }(y,x_{3})\in\Gamma^{\varepsilon}_{+}\cup\Gamma^{\varepsilon}_{-}.$
In order to establish this equivalence result, we will need to resort to the Kirchhoff-Love assumptions (cf., e.g., page 336 in [8]), according to which any point on a normal to the middle surface remains on the normal to the deformed middle surface after the deformation has taken place, and the distance between such a point and the middle surface remains constant before and after the deformation. The latter part of the Kirchhoff-Love assumptions states that it is not possible to compress the linearly elastic shell.
Let us recall that, apart from being judicious and widely accepted, the Kirchhoff-Love assumptions, which are assumptions on the geometrical nature of the shell, constitute one of the pillars that led Koiter to the initial formulation of the homonymous model [29, 30].
For the sake of completeness, let us recall that the other pillar the original formulation of Koiter’s model is based on is an assumption of mechanical nature devised by F. John according to which if the shell thickness is small enough, the state of stress is “approximately” planar and the stresses parallel to the middle surface vary “approximately linearly” across the thickness, at least “away from the lateral face” (cf. [25, 26]).
In order to establish, under the validity of the Kirchhoff-Love assumptions, that the confinement condition considered in this paper is in general stronger than the classical Signorini condition, for modelling purposes we restrict ourselves to considering smooth displacement fields. This choice is judicious, as it is related to the fact that if the norms $\|v^{\varepsilon}_{i}\|_{\mathcal{C}^{1}(\overline{\Omega^{\varepsilon}})}$ are small enough, the mapping $(\bm{\Theta}^{\varepsilon}+v^{\varepsilon}_{i}\bm{g}^{i,\varepsilon})$ is also an injective immersion (cf. section 2).
Lemma 4.2.
Let $\bm{\theta}\in\mathcal{C}^{3}(\overline{\omega};\mathbb{E}^{3})$ be an injective immersion and let $\varepsilon>0$ be small enough so that $\bm{\Theta}^{\varepsilon}\in\mathcal{C}^{3}(\overline{\Omega^{\varepsilon}};\mathbb{E}^{3})$ is also an injective immersion. Let $\bm{v}^{\varepsilon}\in\bm{\mathcal{C}}^{1}(\overline{\Omega^{\varepsilon}})$ be such that $(\bm{\Theta}^{\varepsilon}+v_{i}^{\varepsilon}\bm{g}^{i,\varepsilon})$ is an injective immersion.
Under the validity of the Kirchhoff-Love assumptions, the following conditions are equivalent:
  • (i)
    $\bm{v}^{\varepsilon}\in\bm{U}(\Omega^{\varepsilon})$;
  • (ii)
    $(\bm{\Theta}^{\varepsilon}(y,x_{3}^{\varepsilon})+v_{i}^{\varepsilon}(y,x_{3}^{\varepsilon})\bm{g}^{i,\varepsilon}(y,x_{3}^{\varepsilon}))\cdot\bm{q}\geq 0$, for all $(y,x_{3}^{\varepsilon})\in\Gamma^{\varepsilon}_{+}\cup\Gamma^{\varepsilon}_{-}$.
Proof.
Let us first show that (i) implies (ii). If there was a point $(\hat{y},\hat{x}_{3}^{\varepsilon})\in\Gamma^{\varepsilon}_{+}\cup\Gamma^{\varepsilon}_{-}$ such that
$(\bm{\theta}(\hat{y})+\hat{x}_{3}^{\varepsilon}\bm{a}^{3}(\hat{y})+v_{i}^{\varepsilon}(\hat{y},\hat{x}_{3}^{\varepsilon})\bm{g}^{i,\varepsilon}(\hat{y},\hat{x}_{3}^{\varepsilon}))\cdot\bm{q}<0,$
then we could find a ball centred at the given point $(\hat{y},\hat{x}_{3}^{\varepsilon})$ with small enough radius $\rho>0$, that we denote by $B_{\rho}$, such that:
$(\bm{\theta}(y)+x_{3}^{\varepsilon}\bm{a}^{3}(y)+v_{i}^{\varepsilon}(y,x_{3}^{\varepsilon})\bm{g}^{i,\varepsilon}(y,x_{3}^{\varepsilon}))\cdot\bm{q}<0,$
for all $(y,x_{3}^{\varepsilon})\in\overline{\Omega^{\varepsilon}}\cap B_{\rho}$. The latter contradicts (i), namely, the fact that $\bm{v}^{\varepsilon}\in\bm{U}(\Omega^{\varepsilon})$. Therefore, it must occur that (ii) holds.
Let us now show the converse implication. In light of the assumed Kirchhoff-Love assumptions we have that, for each $y\in\overline{\omega}$, if the closed segment $\{\bm{\Theta}^{\varepsilon}(y,x_{3}^{\varepsilon}),-\varepsilon\leq x_{3}^{\varepsilon}\leq\varepsilon\}$ is orthogonal to the plane that is tangent to $\bm{\theta}(\overline{\omega})$ at $\bm{\theta}(y)$ then the segment
$s_{y}^{\varepsilon}(r):=(1-r)\left(\bm{\theta}(y)-\varepsilon\bm{a}^{3}(y)+v_{i}^{\varepsilon}(y,-\varepsilon)\bm{g}^{i,\varepsilon}(y,-\varepsilon)\right)+r\left(\bm{\theta}(y)+\varepsilon\bm{a}^{3}(y)+v_{i}^{\varepsilon}(y,\varepsilon)\bm{g}^{i,\varepsilon}(y,\varepsilon)\right),$
where $0\leq r\leq 1$ is also orthogonal to the tangent plane to the deformed middle surface at the point $(\bm{\theta}(y)+v_{i}^{\varepsilon}(y,0)\bm{g}^{i,\varepsilon}(y,0))$.
Recalling that we restrict our attention to displacements for which the corresponding deformation is an injective immersion, we have that for all $y_{1},y_{2}\in\overline{\omega}$, with $y_{1}\neq y_{2}$,
$\left(\bigcup_{0\leq r\leq 1}s_{y_{1}}^{\varepsilon}(r)\right)\cap\left(\bigcup_{0\leq r\leq 1}s_{y_{2}}^{\varepsilon}(r)\right)=\emptyset,$
i.e., the segments corresponding to distinct points in $\overline{\omega}$ do not intersect. Therefore, each point in the deformed reference configuration of the shell belongs to one and only one segment orthogonal to the middle surface of the deformed shell.
Since (ii) is assumed to hold, we have that:
$\left[(1-r)\left(\bm{\theta}(y)-\varepsilon\bm{a}^{3}(y)+v_{i}^{\varepsilon}(y,-\varepsilon)\bm{g}^{i,\varepsilon}(y,-\varepsilon)\right)+r\left(\bm{\theta}(y)+\varepsilon\bm{a}^{3}(y)+v_{i}^{\varepsilon}(y,\varepsilon)\bm{g}^{i,\varepsilon}(y,\varepsilon)\right)\right]\cdot\bm{q}\geq 0,$
for all $0\leq r\leq 1$, for all $y\in\overline{\omega}$.
By the arbitrariness of $y\in\overline{\omega}$, we infer that $\bm{v}^{\varepsilon}\in\bm{U}(\Omega^{\varepsilon})$ and the proof is complete. ∎
Note that since $\Gamma^{\varepsilon}_{+}\cup\Gamma^{\varepsilon}_{-}$ clearly contains the portion of the boundary of the three-dimensional elastic shell that engages contact with the obstacle, we have shown that the confinement condition considered in this paper implies the validity of the classical Signorini condition.
When one considers a family of linearly elastic generalised membrane shells (not specifically of the first kind) whose thickness $2\varepsilon$ approaches zero, a specific Korn’s inequality in curvilinear coordinates (cf., e.g., Theorem 4.1 of [18] or Theorem 5.3-1 of [8]) holds over the fixed domain $\Omega=\omega\times(-1,1)$, according to the following theorem. That the constant $C_{1}$ that appears in this inequality is independent of $\varepsilon>0$ plays a key role in the asymptotic analysis of such a family (see part (i) of the proof of Theorem 5.1). A similar inequality of Korn’s type was established by Kohn & Vogelius [28] in the case where the integration domain is $\Omega^{\varepsilon}$.
Theorem 4.2.
Let $\bm{\theta}\in\mathcal{C}^{3}(\overline{\omega};\mathbb{R}^{3})$ be an immersion. Let there be given a family of linearly elastic flexural shells with the same middle surface $\bm{\theta}(\overline{\omega})$ and thickness $02\varepsilon>$ 0. Define the space
$\bm{V}(\Omega):=\{\bm{v}=(v_{i})\in\bm{H}^{1}(\Omega);\;\bm{v}=\bm{0}\textup{ on }\gamma_{0}\times[-1,1]\}.$
Then there exist constants $\varepsilon_{0}>0$ and $C_{1}>0$ such that
$\left\{\sum_{i}\left\|v_{i}\right\|^{2}_{H^{1}(\Omega)}\right\}^{1/2}\leq\dfrac{C_{1}}{\varepsilon}\left\{\sum_{i,j}\left\|e_{i\|j}(\varepsilon;\bm{v})\right\|^{2}_{L^{2}(\Omega)}\right\}^{1/2}$
for all $0<\varepsilon\leq\varepsilon_{0}$ and all $\bm{v}=(v_{i})\in\bm{V}(\Omega)$. ∎
In order to derive the a priori estimates for the family $\{\bm{u}(\varepsilon)\}_{\varepsilon>0}$ of solutions for Problem $\mathcal{P}(\Omega^{\varepsilon})$, we need to specialise the applied body forces in as way that they are admissible. By this we mean that, for a given $\varepsilon>0$, one can find functions $F^{ij}(\varepsilon)=F^{ji}(\varepsilon)\in L^{2}(\Omega)$ such that
(1) $\int_{\Omega}f^{i}v_{i}\sqrt{g(\varepsilon)}\,\mathrm{d}x=\int_{\Omega}F^{ij}(\varepsilon)e_{i\|j}(\varepsilon;\bm{v})\sqrt{g(\varepsilon)}\,\mathrm{d}x,\quad\textup{ for all }\bm{v}=(v_{i})\in\bm{V}(\Omega)\textup{ andall }0<\varepsilon\leq\varepsilon_{0},$
$F^{ij}(\varepsilon)\to F^{ij},\quad\textup{ in }L^{2}(\Omega)\textup{ as }\varepsilon\to 0.$
The details leading to this specialisation for the applied body forces can be found in, for instance, Section 5.5 of [8]. If the applied body forces are admissible, we have that there exists a constant $\kappa_{0}>0$ independent of $\varepsilon$ such that
(2) $\left|\int_{\Omega}F^{ij}(\varepsilon)e_{i\|j}(\varepsilon;\bm{v})\sqrt{g(\varepsilon)}\,\mathrm{d}x\right|\leq\kappa_{0}\left\{\sum_{i,j}\left\|e_{i\|j}(\varepsilon;\bm{v})\right\|_{L^{2}(\Omega)}^{2}\right\}^{1/2},$
for all $0<\varepsilon<\varepsilon_{0}$ and for all $\bm{v}\in\bm{V}(\Omega)$.
The difficulty in the study of linearly elastic generalised membranes lies in the fact that the asymptotic analysis as $\varepsilon\to 0$ involves, at least in the obstacle free case, certain abstract completions. More precisely, for linearly elastic generalised membrane shells of the first kind, the results of the asymptotic analysis are posed (cf., e.g., Theorem 5.6-1 of [8]) over the space
$\bm{V}_{M}^{\sharp}(\omega):=\overline{\bm{V}(\omega)}^{|\cdot|_{\omega}^{M}},$
whereas, for linearly elastic generalised membrane shells of the second kind, the results of the asymptotic analysis are posed (cf., e.g., Theorem 5.6-2 of [8]) over the space
$\dot{\bm{V}}_{M}^{\sharp}(\omega):=\overline{\bm{V}(\omega)/\bm{V}_{0}(\omega)}^{|\cdot|_{\omega}^{M}}.$
In what follows, given a function $v\in L^{2}(\Omega)$, we denote by $\bar{v}$ the average of $v$ with respect to the transverse variable, i.e.,
$\bar{v}:=\dfrac{1}{2}\int_{-1}^{1}v(\cdot,x_{3})\,\mathrm{d}x_{3},$
and we recall that the properties for averages with respect to the transverse variable are established in Section 4.2 of [8]. In the same spirit, we denote the average with respect to the transverse variable of a vector field $\bm{v}=(v_{i})\in\bm{L}^{2}(\Omega)$ by $\bar{\bm{v}}$, and it is clearly defined component-wise by $\bar{\bm{v}}:=\left(\bar{v}_{i}\right)$.
To carry out the asymptotic analysis for a family of linearly elastic generalised membrane shells, we shall need to resort to the space
$\bm{V}_{0}(\Omega):=\{\bm{v}\in\bm{H}^{1}(\Omega);\bm{v}=\bm{0}\textup{ on }\Gamma_{0},\partial_{3}\bm{v}=\bm{0}\textup{ in }\Omega,\textup{ and }\gamma_{\alpha\beta}(\bar{\bm{v}})=0\textup{ in }\omega\},$
which is the “three-dimensional analogue” of the space $\bm{V}_{0}(\omega)$ introduced beforehand. We shall also need to resort to the semi-norm $|\cdot|_{\Omega}^{M}$ defined by:
$|\bm{v}|_{\Omega}^{M}:=\left\{\|\partial_{3}\bm{v}\|_{L^{2}(\omega)}^{2}+\left(|\bar{\bm{v}}|_{\omega}^{M}\right)^{2}\right\}^{1/2},\quad\textup{ for all }\bm{v}\in\bm{V}(\Omega).$
We note in passing that $\bm{V}_{0}(\omega)=\{\bm{0}\}$ if and only if $\bm{V}_{0}(\Omega)=\{\bm{0}\}$, or, equivalently, that $|\cdot|_{\Omega}^{M}$ is a norm over the space $\bm{V}(\Omega)$ if and only if $|\cdot|_{\omega}^{M}$ is a norm over the space $\bm{V}(\omega)$. In other words, we can also say that a linearly elastic generalised membrane shell is of the first kind if and only if $\bm{V}_{0}(\Omega)=\{\bm{0}\}$.
5. Rigorous asymptotic analysis for a family of linearly elastic generalised membranes of the first kind
In this section and in the next one, we restrict our attention to linearly elastic generalised membrane shells of the first kind.
This section is devoted to showing, in the same spirit as [15] (see also Theorem 5.6-1 of [8]), that the solutions $\bm{u}(\varepsilon)$ of the (scaled) three-dimensional problems $\mathcal{P}(\varepsilon;\Omega)$ converge - as $\varepsilon$ approaches zero - to the solution of a two-dimensional limit variational problem, denoted by $\mathcal{P}_{M}^{\sharp}(\omega)$ in what follows.
We will show that the solutions $\bm{u}(\varepsilon)$ for Problem $\mathcal{P}(\varepsilon;\Omega)$ converge in the completion of the space $\bm{V}(\Omega)$ with respect to the norm $|\cdot|_{\Omega}^{M}$. As a result, the two-dimensional limit model we shall recover as a result of the completion of the forthcoming asymptotic analysis will be an abstract variational inequality for which the competitors to the role of solution will vary in the abstract completion of a non-empty, closed, and convex set with respect to a certain norm.
We denote by $B_{M}^{\sharp}(\cdot,\cdot):\bm{V}_{M}^{\sharp}(\omega)\times\bm{V}_{M}^{\sharp}(\omega)\to\mathbb{R}$ the unique continuous extension to the space $\bm{V}_{M}^{\sharp}(\omega)\times\bm{V}_{M}^{\sharp}(\omega)$ of the bilinear form characterising the energy for linearly elastic elliptic membrane shells, which is denoted and defined by:
$(\bm{\eta},\bm{\xi})\in\bm{V}(\omega)\times\bm{V}(\omega)\mapsto B_{M}(\bm{\eta},\bm{\xi}):=\int_{\omega}a^{\alpha\beta\sigma\tau}\gamma_{\sigma\tau}(\bm{\eta})\gamma_{\alpha\beta}(\bm{\xi})\sqrt{a}\,\mathrm{d}y.$
For each $\varepsilon>0$, the linear mapping $L(\varepsilon):\bm{V}(\Omega)\to\mathbb{R}$ is defined by:
$L(\varepsilon)(\bm{v}):=\int_{\Omega}f^{i}v_{i}\sqrt{g(\varepsilon)}\,\mathrm{d}x,\quad\textup{ for all }\bm{v}=(v_{i})\in\bm{V}(\Omega).$
The linear mapping $L_{M}^{\sharp}:\bm{V}_{M}^{\sharp}(\omega)\to\mathbb{R}$, whose rigorous definition will be introduced in the proof of the next theorem, reflects the asymptotic behaviour of the admissible applied body forces $F^{ij}(\varepsilon)$ introduced beforehand. Define the set
$\bm{U}(\omega):=\{\bm{\eta}=(\eta_{i})\in\bm{H}^{1}(\omega);\eta_{i}=0\textup{on }\gamma_{0}\textup{ and }(\bm{\theta}(y)+\eta_{i}(y)\bm{a}^{i}(y))\cdot\bm{q}\geq 0\textup{ for a.a. }y\in\omega\},$
and define the corresponding abstract completion with respect to the norm $|\cdot|_{\omega}^{M}$ as follows:
$\bm{U}_{M}^{\sharp}(\omega):=\overline{\bm{U}(\omega)}^{|\cdot|_{\omega}^{M}}.$
We are thus in a position to state the two-dimensional limit problem, that we denote by $\mathcal{P}_{M}^{\sharp}(\omega)$, as follows:
Problem $\mathcal{P}_{M}^{\sharp}(\omega)$.
Find $\bm{\zeta}\in\bm{U}_{M}^{\sharp}(\omega)$ that satisfies the following variational inequalities:
$B_{M}^{\sharp}(\bm{\zeta},\bm{\eta}-\bm{\zeta})\geq L_{M}^{\sharp}(\bm{\eta}-\bm{\zeta}),\quad\textup{ for all }\bm{\eta}\in\bm{U}_{M}^{\sharp}(\omega).$
$\blacksquare$
Differently from the cases where the linearly elastic shell under consideration is a linearly elastic elliptic membrane shells [19, 20] or a linearly elastic flexural shell [41], the existence and uniqueness of solutions for Problem $\mathcal{P}_{M}^{\sharp}(\omega)$ is not obvious, as it results on a preliminary estimate that we will establish in the forthcoming theorem, which constitutes the first new result contained in this paper.
Prior to establishing this new result, we need to prove a preparatory lemma. The proof of the preparatory lemma we want to establish hinges on the Kirchhoff-Love assumptions, thanks to which it was possible to bridge the confinement condition considered in this paper with the classical Signorini condition (cf. Lemma 4.2). Since we are in the range of infinitesimal displacements, it is licit to assume that if:
(3) $\min_{y\in\overline{\omega}}(\bm{a}^{3}(y)\cdot\bm{q})>0,$
then there exists a constant $M>0$ independent of the choice of the admissible displacement $\bm{\eta}$ for which the following relation holds as well
(4) $\bm{a}_{3}(\bm{\eta})(y)\cdot\bm{q}\geq M,\quad\textup{ for a.a. }y\in\omega,$
where $\bm{a}_{3}(\bm{\eta})$ is the usual unit vector normal to the deformed middle surface corresponding to the displacement occurring at the middle surface incurs. We recall that the relation (3) is one of the sufficient conditions ensuring the “density property” formulated in [19, 20]. We also observe that the condition (4) is implied by the fact that $\overline{\bm{u}(\varepsilon)}\in\bm{H}^{3}(\omega)$ in view of the Rellich-Kondrašov Theorem (cf., e.g., Theorem 6.6-3 in [10]). However, in general, the latter higher regularity does not hold [54], and we thus settle for assuming (4) in view of the infinitesimality of the admissible displacement fields magnitude.
Lemma 5.1.
Let $\bm{u}(\varepsilon)\in\bm{U}(\Omega)$ be the unique solution of Problem $\mathcal{P}(\varepsilon;\Omega)$. Assume that:
$\min_{y\in\overline{\omega}}(\bm{a}^{3}(y)\cdot\bm{q})>0.$
Then, we have that $\overline{\bm{u}(\varepsilon)}\in\bm{U}(\omega)$.
Proof.
Since the Kirchhoff-Love assumptions are valid, we have that the following formula holds (cf., e.g., page 372 in [8]):
$u_{i}(\varepsilon)\bm{g}^{i}(\varepsilon)=u_{i}(\varepsilon)(\cdot,0)\bm{a}^{i}+\varepsilon x_{3}\{\bm{a}_{3}(\bm{u}(\varepsilon)(\cdot,0))-\bm{a}_{3}\},\quad\textup{ a.e. in }\Omega.$
The properties of the covariant and contravariant bases (cf., e.g., Lemma 4.1 and [9]) give at once that:
$u_{\beta}(\varepsilon)$ $=u_{\beta}(\varepsilon)(\cdot,0)-\varepsilon x_{3}u_{i}(\varepsilon)(\cdot,0)b_{\beta\tau}a^{i\tau}+\varepsilon x_{3}\bm{a}_{3}(\bm{u}(\varepsilon)(\cdot,0))\cdot\bm{a}_{\beta}$
$\quad-\varepsilon^{2}x_{3}^{2}b_{\beta\tau}\bm{a}_{3}(\bm{u}(\varepsilon)(\cdot,0))\cdot\bm{a}^{\tau},$
$u_{3}(\varepsilon)$ $=u_{3}(\varepsilon)(\cdot,0)+\varepsilon x_{3}\bm{a}_{3}(\bm{u}(\varepsilon)(\cdot,0))\cdot\bm{a}_{3}-\varepsilon x_{3}.$
Therefore, we have that the average of $\bm{u}(\varepsilon)$ satisfies the following relations:
(5) $\overline{u_{\beta}(\varepsilon)}$ $=u_{\beta}(\varepsilon)(\cdot,0)-\dfrac{2}{3}\varepsilon^{2}b_{\beta\tau}\bm{a}_{3}(\bm{u}(\varepsilon)(\cdot,0))\cdot\bm{a}^{\tau},$
$\overline{u_{3}(\varepsilon)}$ $=u_{3}(\varepsilon)(\cdot,0).$
In view of (5), we have that for a.a. $y\in\omega$ the following relations hold:
$\left(\bm{\theta}(y)+\overline{u_{i}(\varepsilon)}(y)\bm{a}^{i}(y)\right)\cdot\bm{q}=\left(\bm{\theta}(y)+u_{i}(\varepsilon)(y,0)\bm{a}^{i}\right)\cdot\bm{q}-\dfrac{2}{3}\varepsilon^{2}b_{\beta\tau}(\bm{a}_{3}(\bm{u}(\varepsilon)(y,0))\cdot\bm{a}^{\tau})\bm{a}^{\beta}\cdot\bm{q}$
$=\left(\bm{\theta}(y)+u_{i}(\varepsilon)(y,0)\bm{a}^{i}-\varepsilon\bm{a}_{3}(\bm{u}(\varepsilon)(y,0))\right)\cdot\bm{q}+\varepsilon\bm{a}_{3}(\bm{u}(\varepsilon)(y,0))\cdot\bm{q}-\dfrac{2}{3}\varepsilon^{2}b_{\beta\tau}(\bm{a}_{3}(\bm{u}(\varepsilon)(y,0))\cdot\bm{a}^{\tau})(\bm{a}^{\beta}\cdot\bm{q}).$
In light of the Kirchhoff-Love assumptions, it is possible to observe that the first addend on the right-hand side of the last inequality represents the deformation of the lower face of the shell at the given point $y\in\omega$, since all the points along a normal to the middle surface line on a normal to the deformed middle surface. Therefore, by Lemma 4.2, we have that:
$\left(\bm{\theta}(y)+u_{i}(\varepsilon)(y,0)\bm{a}^{i}-\varepsilon\bm{a}_{3}(\bm{u}(\varepsilon)(y,0))\right)\cdot\bm{q}\geq 0,\quad\textup{ for a.a. }y\in\omega.$
In light of (4), we have that:
$\bm{a}_{3}(\bm{u}(\varepsilon)(y,0))\cdot\bm{q}\geq M,\quad\textup{ for a.a. }y\in\omega,$
where the constant $M$ is independent of the displacement and, therefore, it is independent of $\varepsilon$.
Therefore, if we choose $\varepsilon>0$ sufficiently small, we have that the smoothness of $b_{\beta\tau}$, the smoothness of $\bm{a}^{\tau}$, and the fact that $\bm{a}_{3}(\bm{u}(\varepsilon)(y,0))$ has unit norm (cf., e.g., formula (1.3.4) in [1]) give that
$\left(\bm{\theta}(y)+\overline{u_{i}(\varepsilon)}(y)\bm{a}^{i}(y)\right)\cdot\bm{q}\geq 0,\quad\textup{ for a.a. }y\in\omega,$
as it was to be proved. ∎
We are now ready to prove the first main result of this paper.
Theorem 5.1.
Assume that $\bm{\theta}\in\mathcal{C}^{3}(\overline{\omega};\mathbb{E}^{3})$ is an injective immersion. Consider a family of linearly elastic generalised membrane shells of the first kind, each of which having the same middle surface $\bm{\theta}(\overline{\omega})$, and thickness $2\varepsilon$ approaching zero. Assume that each of the linearly elastic generalised membrane shells of the first kind under consideration is subjected to a boundary condition of place along a portion of its lateral face having $\bm{\theta}(\gamma_{0})$ as its middle curve, and that each of these shells is subjected to applied body forces that are admissible (cf. section 4).
Assume that (3) holds, namely:
$\min_{y\in\overline{\omega}}(\bm{a}^{3}(y)\cdot\bm{q})>0.$
For each $0<\varepsilon\leq\varepsilon_{0}$, let $\bm{u}(\varepsilon)\in\bm{U}(\Omega)$ denote the unique solution of Problem $\mathcal{P}(\varepsilon;\Omega)$. Define the space
$\bm{V}_{M}^{\sharp}(\Omega):=\overline{\bm{V}(\Omega)}^{|\cdot|_{\Omega}^{M}},$
and define the set
$\bm{U}_{M}^{\sharp}(\omega):=\overline{\bm{U}(\omega)}^{|\cdot|_{\omega}^{M}}\subset\bm{V}_{M}^{\sharp}(\omega).$
Then there exists $\bm{u}\in\bm{V}_{M}^{\sharp}(\Omega)$ and $\bm{\zeta}\in\bm{U}_{M}^{\sharp}(\omega)$ such that:
$\bm{u}(\varepsilon)$ $\to\bm{u},\quad\textup{ in }\bm{V}_{M}^{\sharp}(\Omega)\textup{ as }\varepsilon\to 0,$
$\overline{\bm{u}(\varepsilon)}$ $\to\bm{\zeta},\quad\textup{ in }\bm{V}_{M}^{\sharp}(\omega)\textup{ as }\varepsilon\to 0.$
Define
$a^{\alpha\beta\sigma\tau}$ $:=\dfrac{4\lambda\mu}{\lambda+2\mu}a^{\alpha\beta}a^{\sigma\tau}+2\mu(a^{\alpha\sigma}a^{\beta\tau}+a^{\alpha\tau}a^{\beta\sigma}),$
$\varphi^{\alpha\beta}$ $:=\int_{-1}^{1}\left\{F^{\alpha\beta}-\dfrac{\lambda}{\lambda+2\mu}a^{\alpha\beta}F^{33}\right\}\,\mathrm{d}x_{3}\in L^{2}(\omega),$
$L_{M}(\bm{\eta})$ $:=\int_{\omega}\varphi^{\alpha\beta}\gamma_{\alpha\beta}(\bm{\eta})\sqrt{a}\,\mathrm{d}y,\quad\textup{ for all }\bm{\eta}\in\bm{V}(\omega),$
where the functions $F^{ij}\in L^{2}(\Omega)$ are those entering the definition of admissible forces, and $B_{M}^{\sharp}$ and $L_{M}^{\sharp}$ denote the unique continuous extensions from $\bm{V}(\omega)$ to $\bm{V}_{M}^{\sharp}(\omega)$ of the bilinear form $B_{M}(\cdot,\cdot)$ and of the linear form $L_{M}$, respectively. Then, the limit $\bm{\zeta}$ belongs to the set $\bm{U}_{M}^{\sharp}(\omega)$, and it is the unique solution of Problem $\mathcal{P}_{M}^{\sharp}(\omega)$.
Proof.
In what follows, we resort to the following compact notation:
$e_{i\|j}(\varepsilon):=e_{i\|j}(\varepsilon;\bm{u}(\varepsilon)).$
The outline of the proof, which is broken into ten parts numbered (i)–(x), is inspired by the proof outline of Theorem 5.6-1 of [8] (itself adapted from the original result by Ciarlet & Lods contained in [15]), where no confinement conditions are imposed. We will see that addressing the confinement condition we are taking into account will require extra care in some parts of the asymptotic analysis.
(i) There exists $0<\varepsilon_{1}\leq\varepsilon_{0}$, and there exists a constant $c_{0}>0$ independent of $\varepsilon$ such that, for all $0<\varepsilon\leq\varepsilon_{1}$, we have that:
$|\bm{v}|_{\Omega}^{M}\leq c_{0}\left\{\sum_{i,j}\|e_{i\|j}(\varepsilon;\bm{v})\|_{L^{2}(\Omega)}^{2}\right\}^{1/2},\quad\textup{ for all }\bm{v}\in\bm{V}(\Omega).$
The proof is the same as part (i) in Theorem 5.6-1 of [8] and is, for this reason, omitted.
(ii) The semi-norms $|\bm{u}(\varepsilon)|_{\Omega}^{M}$, $\left|\overline{\bm{u}(\varepsilon)}\right|_{\omega}^{M}$, and the norms $\|\varepsilon\bm{u}(\varepsilon)\|_{\bm{H}^{1}(\Omega)}$ and $\|e_{i\|j}(\varepsilon)\|_{L^{2}(\Omega)}$ are bounded independently of $0<\varepsilon\leq\varepsilon_{1}$. Therefore, up to passing to a subsequence, there exist $\bm{u}\in\bm{V}_{M}^{\sharp}(\Omega)$, $\bm{u}^{-1}=(u_{i}^{(-1)})\in\bm{V}(\Omega)$, $e_{i\|j}\in L^{2}(\Omega)$, and $\bm{\zeta}\in\bm{U}_{M}^{\sharp}(\omega)$ for which the following convergences take place as $\varepsilon\to 0$:
$\bm{u}(\varepsilon)$ $\rightharpoonup\bm{u},\quad\textup{ in }\bm{V}_{M}^{\sharp}(\Omega),$
$\varepsilon\bm{u}(\varepsilon)$ $\rightharpoonup\bm{u}^{(-1)},\quad\textup{ in }\bm{H}^{1}(\Omega),$
$e_{i\|j}(\varepsilon)$ $\rightharpoonup e_{i\|j},\quad\textup{ in }\bm{L}^{2}(\Omega),$
$\partial_{3}u_{3}(\varepsilon)$ $=\varepsilon e_{3\|3}(\varepsilon)\to 0,\quad\textup{ in }L^{2}(\Omega),$
$\overline{\bm{u}(\varepsilon)}$ $\rightharpoonup\bm{\zeta},\quad\textup{ in }\bm{U}_{M}^{\sharp}(\omega).$
Specialising $\bm{v}=\bm{0}$ in the variational inequalities of Problem $\mathcal{P}(\varepsilon;\Omega)$, and recalling that the applied body forces are admissible (see (1) and (2)), we obtain, as a result of an application of part (i), that
$\dfrac{\sqrt{g_{0}}}{C_{e}c_{0}^{2}}\left(|\bm{u}(\varepsilon)|_{\Omega}^{M}\right)^{2}\leq\dfrac{\sqrt{g_{0}}}{C_{e}}\sum_{i,j}\|e_{i\|j}(\varepsilon)\|_{L^{2}(\Omega)}^{2}\leq\int_{\Omega}A^{ijk\ell}(\varepsilon)e_{k\|\ell}(\varepsilon)e_{i\|j}(\varepsilon)\sqrt{g(\varepsilon)}\,\mathrm{d}x$
$\leq\int_{\Omega}F^{ij}(\varepsilon)e_{i\|j}(\varepsilon)\sqrt{g(\varepsilon)}\,\mathrm{d}x\leq\kappa_{0}\sqrt{g_{1}}\left\{\sum_{i,j}\|e_{i\|j}(\varepsilon)\|_{L^{2}(\Omega)}^{2}\right\}^{1/2},$
for some $\kappa_{0}>0$ independent of $\varepsilon$. As a result, we obtain that
[…]
An application of Theorem 4.2-1 in [8] gives that
$\left|\overline{\bm{u}(\varepsilon)}\right|_{\omega}^{M}\leq|\bm{u}(\varepsilon)|_{\Omega}^{M},$
and, an application of Theorem 4.2 gives the boundedness of the sequence $\{\varepsilon\|\bm{u}(\varepsilon)\|_{\bm{H}^{1}(\Omega)}\}_{\varepsilon>0}$ independently of $0<\varepsilon\leq\varepsilon_{1}$.
In order to show that $\bm{\zeta}\in\bm{U}_{M}^{\sharp}(\omega)$, it suffices to verify that $\bm{U}_{M}^{\sharp}(\omega)$ is convex (note that $\bm{U}_{M}^{\sharp}(\omega)$ is - by definition - (strongly) closed with respect to the norm $|\cdot|_{\omega}^{M}$), so that a classical result in Functional Analysis ensures that $\bm{U}_{M}^{\sharp}(\omega)$ is also weakly closed (cf., e.g., [2]).
Let $\bm{\zeta}_{1},\bm{\zeta}_{2}\in\bm{U}_{M}^{\sharp}(\omega)$, and let $0\leq r\leq 1$. Consider the sequences $\{\bm{\zeta}_{1}^{(k)}\}_{k\geq 0}$ and $\{\bm{\zeta}_{2}^{(k)}\}_{k\geq 0}$ in $\bm{U}(\omega)$ such that:
$|\bm{\zeta}_{1}^{(k)}-\bm{\zeta}_{1}|_{\omega}^{M}$ $\to 0,\quad\textup{ as }k\to\infty,$
$|\bm{\zeta}_{2}^{(k)}-\bm{\zeta}_{2}|_{\omega}^{M}$ $\to 0,\quad\textup{ as }k\to\infty.$
An application of the triangle inequality gives that:
$\left|\left(r\bm{\zeta}_{1}^{(k)}+(1-r)\bm{\zeta}_{2}^{(k)}\right)-\left(r\bm{\zeta}_{1}+(1-r)\bm{\zeta}_{2}\right)\right|_{\omega}^{M}\leq r|\bm{\zeta}_{1}^{(k)}-\bm{\zeta}_{1}|_{\omega}^{M}+(1-r)|\bm{\zeta}_{2}^{(k)}-\bm{\zeta}_{2}|_{\omega}^{M}\to 0,\quad\textup{ as }k\to\infty.$
Combining the previous convergence and the convexity of the set $\bm{U}(\omega)$ gives that
$\left(r\bm{\zeta}_{1}+(1-r)\bm{\zeta}_{2}\right)\in\bm{U}_{M}^{\sharp}(\omega),$
and the sought convexity is thus obtained, thus showing that $\bm{U}_{M}^{\sharp}(\omega)$ is weakly closed.
As a result, combining the fact that the elements $\overline{\bm{u}(\varepsilon)}\in\bm{U}(\omega)$ (cf. Lemma 5.1) are bounded independently of $0<\varepsilon\leq\varepsilon_{1}$ with respect to the norm $|\cdot|_{\omega}^{M}$ with the fact that $\bm{U}_{M}^{\sharp}(\omega)$ is weakly closed in turn gives that
$\overline{\bm{u}(\varepsilon)}\rightharpoonup\bm{\zeta},\quad\textup{ in }\bm{U}_{M}^{\sharp}(\omega)\textup{ as }\varepsilon\to 0.$
(iii) There exists a constant $0<\varepsilon_{2}\leq\varepsilon_{1}$ such that, given any $0<\varepsilon\leq\varepsilon_{2}$ and any vector field $\bm{\varphi}=(\varphi_{i})\in\bm{\mathcal{D}}(\Omega)$, one can find a vector field $\bm{v}(\varepsilon;\bm{\varphi})=(v_{i}(\varepsilon;\bm{\varphi})):\Omega\to\mathbb{R}^{3}$ with the following properties:
$\bm{v}(\varepsilon;\bm{\varphi})\in\bm{U}(\Omega)\quad\textup{ and }\quad\partial_{3}v_{i}(\varepsilon;\bm{\varphi})=\varphi_{i}\textup{ in }\Omega,\textup{ for all }0<\varepsilon\leq\varepsilon_{2}.$
Moreover, there exists a constant $C(\bm{\varphi})$ such that:
$\|\bm{v}(\varepsilon;\bm{\varphi})\|_{\bm{H}^{1}(\Omega)}\leq C(\bm{\varphi}),\quad\textup{ for all }0<\varepsilon\leq\varepsilon_{2}.$
The proof is identical to the proof of part (iv) in [20] and is, for this reason, omitted.
(iv) The weak limits $e_{i\|j}\in L^{2}(\Omega)$ found in part (ii) take the following forms:
$e_{\alpha\|3}$ $=\dfrac{1}{2\mu}a_{\alpha\beta}F^{\beta 3},$
$e_{3\|3}$ $=-\dfrac{\lambda}{\lambda+2\mu}a^{\alpha\beta}e_{\alpha\|\beta}+\dfrac{F^{33}}{\lambda+2\mu}.$
Let the vector field $\bm{\varphi}=(\varphi_{i})\in\bm{\mathcal{D}}(\Omega)$ be given, for a given $0<\varepsilon\leq\varepsilon_{2}$, let $\bm{v}(\varepsilon;\bm{\varphi})$ be the corresponding vector field defined in part (iii). Specialising $\bm{v}=\bm{v}(\varepsilon;\bm{\varphi})$ in the variational inequalities of Problem $\mathcal{P}(\varepsilon;\Omega)$. Multiplying the aforementioned variational inequalities by $\varepsilon>0$, and recalling that $A^{\alpha\beta\sigma 3}(\varepsilon)=A^{\alpha 333}(\varepsilon)=0$, we obtain that:
(6) $\int_{\Omega}A^{\alpha\beta\sigma\tau}(\varepsilon)e_{\sigma\|\tau}(\varepsilon)\left\{\varepsilon e_{\alpha\|\beta}(\varepsilon;\bm{v}(\varepsilon;\bm{\varphi}))-\varepsilon e_{\alpha\|\beta}(\varepsilon)\right\}\sqrt{g(\varepsilon)}\,\mathrm{d}x$
$\quad+\int_{\Omega}A^{\sigma\tau 33}(\varepsilon)e_{3\|3}(\varepsilon)\left\{\varepsilon e_{\sigma\|\tau}(\varepsilon;\bm{v}(\varepsilon;\bm{\varphi}))-\varepsilon e_{\sigma\|\tau}(\varepsilon)\right\}\sqrt{g(\varepsilon)}\,\mathrm{d}x$
$\quad+4\int_{\Omega}A^{\alpha 3\sigma 3}(\varepsilon)e_{\sigma\|3}(\varepsilon)\left\{\varepsilon e_{\alpha\|3}(\varepsilon;\bm{v}(\varepsilon;\bm{\varphi}))-\varepsilon e_{\alpha\|3}(\varepsilon)\right\}\sqrt{g(\varepsilon)}\,\mathrm{d}x$
$\quad+\int_{\Omega}A^{33\sigma\tau}(\varepsilon)e_{\sigma\|\tau}(\varepsilon)\left\{\varepsilon e_{3\|3}(\varepsilon;\bm{v}(\varepsilon;\bm{\varphi}))-\varepsilon e_{3\|3}(\varepsilon)\right\}\sqrt{g(\varepsilon)}\,\mathrm{d}x$
$\quad+\int_{\Omega}A^{3333}(\varepsilon)e_{3\|3}(\varepsilon)\left\{\varepsilon e_{3\|3}(\varepsilon;\bm{v}(\varepsilon;\bm{\varphi}))-\varepsilon e_{3\|3}(\varepsilon)\right\}\sqrt{g(\varepsilon)}\,\mathrm{d}x$
$\geq\int_{\Omega}F^{\alpha\beta}(\varepsilon)\left\{\varepsilon e_{\alpha\|\beta}(\varepsilon;\bm{v}(\varepsilon;\bm{\varphi}))-\varepsilon e_{\alpha\|\beta}(\varepsilon)\right\}\sqrt{g(\varepsilon)}\,\mathrm{d}x$
$\quad+2\int_{\Omega}F^{\alpha 3}(\varepsilon)\left\{\varepsilon e_{\alpha\|3}(\varepsilon;\bm{v}(\varepsilon;\bm{\varphi}))-\varepsilon e_{\alpha\|3}(\varepsilon)\right\}\sqrt{g(\varepsilon)}\,\mathrm{d}x$
$\quad+\int_{\Omega}F^{33}(\varepsilon)\left\{\varepsilon e_{3\|3}(\varepsilon;\bm{v}(\varepsilon;\bm{\varphi}))-\varepsilon e_{3\|3}(\varepsilon)\right\}\sqrt{g(\varepsilon)}\,\mathrm{d}x.$
The relations $\partial_{3}v_{i}(\varepsilon;\bm{\varphi})=\varphi_{i}$, the boundedness of the norms $\|\bm{v}(\varepsilon;\bm{\varphi})\|_{\bm{H}^{1}(\Omega)}$ independently of $0<\varepsilon\leq\varepsilon_{2}$, imply that:
(7) $\varepsilon e_{\alpha\|\beta}(\varepsilon;\bm{v}(\varepsilon;\bm{\varphi}))$ $\to 0,\quad\textup{ in }L^{2}(\Omega),$
$\varepsilon e_{\alpha\|3}(\varepsilon;\bm{v}(\varepsilon;\bm{\varphi}))$ $\to\dfrac{\varphi_{\alpha}}{2},\quad\textup{ in }L^{2}(\Omega),$
$\varepsilon e_{\alpha\|\beta}(\varepsilon;\bm{v}(\varepsilon;\bm{\varphi}))$ $=\varphi_{3},\quad\textup{ a.e. in }\Omega\textup{ for all }0<\varepsilon\leq\varepsilon_{2}.$
Letting $\varepsilon\to 0$ in (6), recalling the definition of admissible applied body forces (1), and recalling convergences in part (ii) and in (7), we obtain that:
(8) $\int_{\Omega}2\mu a^{\alpha\sigma}e_{\sigma\|3}\varphi_{\alpha}\sqrt{a}\,\mathrm{d}x+\int_{\Omega}\left(\lambda a^{\sigma\tau}e_{\sigma\|\tau}+(\lambda+2\mu)e_{3\|3}\right)\varphi_{3}\sqrt{a}\,\mathrm{d}x$
$\geq\int_{\Omega}F^{\alpha 3}\varphi_{\alpha}\sqrt{a}\,\mathrm{d}x+\int_{\Omega}F^{33}\varphi_{3}\sqrt{a}\,\mathrm{d}x.$
As the inequality (8) holds for any vector field $\bm{\varphi}=(\varphi_{i})\in\bm{\mathcal{D}}(\omega)$, we obtain that if we specialise one such vector field in a way that $\varphi_{3}=0$ and $\varphi_{\alpha}$ varies in $\mathcal{D}(\omega)$, then:
$2\mu\begin{pmatrix}a^{11}&a^{12}\\ a^{12}&a^{22}\end{pmatrix}\begin{pmatrix}e_{1\|3}\\ e_{2\|3}\end{pmatrix}=\begin{pmatrix}F^{13}\\ F^{23}\end{pmatrix},\quad\textup{ a.e. in }\Omega.$
The invertibility of the symmetric and positive-definite matrix $(a^{\alpha\beta})$ in turn gives that:
$e_{\alpha\|3}=\dfrac{1}{2\mu}a_{\alpha\beta}F^{\beta 3}.$
The relations
$e_{3\|3}=-\dfrac{\lambda}{\lambda+2\mu}a^{\alpha\beta}e_{\alpha\|\beta}+\dfrac{F^{33}}{\lambda+2\mu},$
can be established likewise, by letting $\varphi_{\alpha}=0$ and letting $\varphi_{3}\in\mathcal{D}(\omega)$ vary arbitrarily.
(v) The whole family $\{\bm{u}(\varepsilon)\}_{\varepsilon>0}$ satisfies:
$\left\{\overline{e_{\alpha\|\beta}(\varepsilon)}-\gamma_{\alpha\beta}\left(\overline{\bm{u}(\varepsilon)}\right)\right\}\to 0,\quad\textup{ in }L^{2}(\omega)\textup{ as }\varepsilon\to 0.$
Consequently, the subsequence considered in part (ii) satisfies:
$\gamma_{\alpha\beta}\left(\overline{\bm{u}(\varepsilon)}\right)\rightharpoonup\overline{e_{\alpha\|\beta}},\quad\textup{ in }L^{2}(\omega)\textup{ as }\varepsilon\to 0.$
The proof is identical to the proof of part (iv) in Theorem 5.6-1 of [8] and is, for this reason, omitted.
(vi) The limits $e_{\alpha\|\beta}$ obtained in part (ii) satisfy the following variational inequalities:
$\int_{\omega}a^{\alpha\beta\sigma\tau}\overline{e_{\sigma\|\tau}}(\gamma_{\alpha\beta}(\bm{\eta})-\overline{e_{\alpha\|\beta}})\sqrt{a}\,\mathrm{d}y\geq\int_{\omega}\varphi^{\alpha\beta}(\gamma_{\alpha\beta}(\bm{\eta})-\overline{e_{\alpha\|\beta}})\sqrt{a}\,\mathrm{d}y,\quad\textup{ for all }\bm{\eta}\in\bm{U}(\omega),$
where
$\varphi^{\alpha\beta}:=\int_{-1}^{1}\left\{F^{\alpha\beta}-\dfrac{\lambda}{\lambda+2\mu}a^{\alpha\beta}F^{33}\right\}\,\mathrm{d}x_{3}\in L^{2}(\omega),$
the functions $F^{ij}\in L^{2}(\Omega)$ being those used in the definition of admissible applied body forces (cf. (1) and (2)).
For a given $\bm{\eta}=(\eta_{i})\in\bm{U}(\omega)$, and for each $0<\varepsilon\leq\varepsilon_{2}$, define the vector field $\bm{v}(\varepsilon;\bm{\eta})=(v_{i}(\varepsilon;\bm{\eta}))$ as follows:
$v_{i}(\varepsilon;\bm{\eta}):=\left[(1-\sqrt{\varepsilon})\eta_{j}\bm{a}^{j}\right]\cdot\bm{g}_{i}(\varepsilon).$
Since $\bm{\eta}\in\bm{U}(\omega)$, since the vector fields $\bm{a}^{j}$ are of class $\bm{\mathcal{C}}^{2}(\overline{\omega})$, and since the vector fields $\bm{g}_{i}(\varepsilon)$ are of class $\bm{\mathcal{C}}^{2}(\overline{\Omega})$, we obtain (cf., e.g., Proposition 9.4 of [2]) that $\bm{v}(\varepsilon;\bm{\eta})\in\bm{H}^{1}(\Omega)$. Moreover, the trace of each component $v_{i}(\varepsilon;\bm{\eta})$ vanishes along $\gamma_{0}\times(-1,1)$. The fact that $\bm{\theta}\cdot\bm{q}>0$ in $\overline{\omega}$ implies that, for a.a. $x\in\Omega$,
$\left[\bm{\theta}+\varepsilon x_{3}\bm{a}^{3}+v_{i}(\varepsilon;\bm{\eta})\bm{g}^{i}(\varepsilon)\right]\cdot\bm{q}$
$=\left[\bm{\theta}+\varepsilon x_{3}\bm{a}^{3}+(\eta_{j}\bm{a}^{j}\cdot\bm{g}_{i}(\varepsilon))\bm{g}^{i}(\varepsilon)\right]\cdot\bm{q}$
$\quad-\sqrt{\varepsilon}(\eta_{j}\bm{a}^{j}\cdot\bm{g}_{i}(\varepsilon))(\bm{g}^{i}(\varepsilon)\cdot\bm{q})$
$=(1-\sqrt{\varepsilon})\underbrace{\left(\bm{\theta}+\eta_{j}\bm{a}^{j}\right)\cdot\bm{q}}_{\geq 0}+\sqrt{\varepsilon}(\underbrace{\bm{\theta}\cdot\bm{q}}_{>0}+\underbrace{\sqrt{\varepsilon}x_{3}\bm{a}^{3}\cdot\bm{q}}_{=\mathcal{O}(\sqrt{\varepsilon})}),$
and we observe that the latter quantity is greater or equal than zero provided that $\varepsilon$ is sufficiently small. This implies that $\bm{v}(\varepsilon;\bm{\eta})\in\bm{U}(\varepsilon;\Omega)$. Since
$v_{j}(\varepsilon;\bm{\eta})\bm{g}^{j}(\varepsilon)=\left(\left((1-\sqrt{\varepsilon})\eta_{i}\bm{a}^{i}\right)\cdot\bm{g}_{j}(\varepsilon)\right)\bm{g}^{j}(\varepsilon)=(1-\sqrt{\varepsilon})\eta_{j}\bm{a}^{j},$
we deduce that:
$\|v_{j}(\varepsilon;\bm{\eta})\bm{g}^{j}(\varepsilon)-\eta_{j}\bm{a}^{j}\|_{\bm{H}^{1}(\Omega)}=\sqrt{\varepsilon}\|\eta_{i}\bm{a}^{i}\|_{\bm{H}^{1}(\Omega)}\to 0,\quad\textup{ as }\varepsilon\to 0.$
The latter convergence in turn implies that:
$v_{i}(\varepsilon;\bm{\eta})=\left(v_{j}(\varepsilon;\bm{\eta})\bm{g}^{j}(\varepsilon)\right)\cdot\bm{g}_{i}(\varepsilon)\to(\eta_{j}\bm{a}^{j})\cdot\bm{a}_{i}=\eta_{i},\quad\textup{ in }H^{1}(\Omega)\textup{ as }\varepsilon\to 0.$
Finally, using Lemma 4.1 and noting that
$v_{\alpha}(\varepsilon;\bm{\eta})$ $=\left((1-\sqrt{\varepsilon})\eta_{j}\bm{a}^{j}\right)\cdot\bm{g}_{\alpha}(\varepsilon)=(1-\sqrt{\varepsilon})\left(\eta_{\alpha}-\varepsilon x_{3}b^{\sigma}_{\alpha}\eta_{\sigma}\right),$
$v_{3}(\varepsilon;\bm{\eta})$ $=\left((1-\sqrt{\varepsilon})\eta_{j}\bm{a}^{j}\right)\cdot\bm{g}_{3}(\varepsilon)=(1-\sqrt{\varepsilon})\eta_{3},$
we deduce, from the definition of the scaled strains $e_{i\|j}(\varepsilon;\bm{v})$, that
(9) $e_{\alpha\|\beta}(\varepsilon;\bm{v}(\varepsilon;\bm{\eta}))$ $=(1-\sqrt{\varepsilon})\left(\gamma_{\alpha\beta}(\bm{\eta})+\varepsilon x_{3}\mathcal{R}_{\alpha\beta}(\varepsilon;\bm{\eta})\right)\to\gamma_{\alpha\beta}(\bm{\eta}),\quad\textup{ in }L^{2}(\Omega)\textup{ as }\varepsilon\to 0$
$e_{\alpha\|3}(\varepsilon;\bm{v}(\varepsilon;\bm{\eta}))$ $=(1-\sqrt{\varepsilon})\left(\frac{1}{2}\left(b^{\sigma}_{\alpha}\eta_{\sigma}+\partial_{\alpha}\eta_{3}\right)+\varepsilon x_{3}\mathcal{R}_{\alpha 3}(\varepsilon;\bm{\eta})\right),$
$e_{3\|3}(\varepsilon;\bm{v}(\varepsilon;\bm{\eta}))$ $=0,$
for some functions $\mathcal{R}_{\alpha j}(\varepsilon;\bm{\eta})\in L^{2}(\Omega)$ that are bounded in $L^{2}(\Omega)$ by a constant independent of $\varepsilon$.
Given $\bm{\eta}\in\bm{U}(\omega)$, specialising the test functions $\bm{v}$ in the variational inequalities by taking $\bm{v}(\varepsilon;\bm{\eta})$, we have that:
(10) $0\leq$ $\int_{\Omega}A^{ijk\ell}(\varepsilon)e_{k\|\ell}(\varepsilon)e_{i\|j}(\varepsilon)\sqrt{g(\varepsilon)}\sqrt{g(\varepsilon)}\,\mathrm{d}x$
$\quad-2\int_{\Omega}A^{ijk\ell}(\varepsilon)e_{k\|\ell}(\varepsilon)e_{i\|j}\sqrt{g(\varepsilon)}\,\mathrm{d}x$
$\quad+\int_{\Omega}A^{ijk\ell}(\varepsilon)e_{k\|\ell}e_{i\|j}\sqrt{g(\varepsilon)}\,\mathrm{d}x$
$\leq\int_{\Omega}A^{ijk\ell}(\varepsilon)e_{k\|\ell}(\varepsilon)e_{i\|j}(\varepsilon;\bm{v}(\varepsilon;\bm{\eta}))\sqrt{g(\varepsilon)}\sqrt{g(\varepsilon)}\,\mathrm{d}x$
$\quad-2\int_{\Omega}A^{ijk\ell}(\varepsilon)e_{k\|\ell}(\varepsilon)e_{i\|j}\sqrt{g(\varepsilon)}\,\mathrm{d}x$
$\quad+\int_{\Omega}A^{ijk\ell}(\varepsilon)e_{k\|\ell}e_{i\|j}\sqrt{g(\varepsilon)}\,\mathrm{d}x$
$\quad-\int_{\Omega}F^{ij}(\varepsilon)\left(e_{i\|j}(\varepsilon;\bm{v}(\varepsilon;\bm{\eta}))-e_{i\|j}(\varepsilon)\right)\sqrt{g(\varepsilon)}\,\mathrm{d}x$
$=\int_{\Omega}A^{\alpha\beta\sigma\tau}(\varepsilon)e_{\sigma\|\tau}(\varepsilon)e_{\alpha\|\beta}(\varepsilon;\bm{v}(\varepsilon;\bm{\eta}))\sqrt{g(\varepsilon)}\,\mathrm{d}x$
$\quad+\int_{\Omega}A^{\sigma\tau 33}(\varepsilon)e_{\sigma\|\tau}(\varepsilon)e_{3\|3}(\varepsilon;\bm{v}(\varepsilon;\bm{\eta}))\sqrt{g(\varepsilon)}\,\mathrm{d}x$
$\quad+4\int_{\Omega}A^{\alpha 3\sigma 3}(\varepsilon)e_{\sigma\|3}(\varepsilon)e_{\alpha\|3}(\varepsilon;\bm{v}(\varepsilon;\bm{\eta}))\sqrt{g(\varepsilon)}\,\mathrm{d}x$
$\quad+\int_{\Omega}A^{33\sigma\tau}(\varepsilon)e_{3\|3}(\varepsilon)e_{\sigma\|\tau}(\varepsilon;\bm{v}(\varepsilon;\bm{\eta}))\sqrt{g(\varepsilon)}\,\mathrm{d}x$
$\quad+\int_{\Omega}A^{3333}(\varepsilon)e_{3\|3}(\varepsilon)e_{3\|3}(\varepsilon;\bm{v}(\varepsilon;\bm{\eta}))\sqrt{g(\varepsilon)}\,\mathrm{d}x$
$\quad-2\int_{\Omega}A^{ijk\ell}(\varepsilon)e_{k\|\ell}(\varepsilon)e_{i\|j}\sqrt{g(\varepsilon)}\,\mathrm{d}x$
$\quad+\int_{\Omega}A^{ijk\ell}(\varepsilon)e_{k\|\ell}e_{i\|j}\sqrt{g(\varepsilon)}\,\mathrm{d}x$
$\quad-\int_{\Omega}F^{\alpha\beta}(\varepsilon)\left(e_{\alpha\|\beta}(\varepsilon;\bm{v}(\varepsilon;\bm{\eta}))-e_{\alpha\|\beta}(\varepsilon)\right)\sqrt{g(\varepsilon)}\,\mathrm{d}x$
$\quad-2\int_{\Omega}F^{\alpha 3}(\varepsilon)\left(e_{\alpha\|3}(\varepsilon;\bm{v}(\varepsilon;\bm{\eta}))-e_{\alpha\|3}(\varepsilon)\right)\sqrt{g(\varepsilon)}\,\mathrm{d}x$
$\quad-\int_{\Omega}F^{33}(\varepsilon)\left(e_{3\|3}(\varepsilon;\bm{v}(\varepsilon;\bm{\eta}))-e_{3\|3}(\varepsilon)\right)\sqrt{g(\varepsilon)}\,\mathrm{d}x.$
The assumed admissibility of the applied body forces (viz. (1) and (2)), the weak convergences in part (ii), the relationships satisfied by the weak limits $e_{i\|j}$, and the asymptotic behaviour of the vector fields $\bm{v}(\varepsilon;\bm{\eta})$ as $\varepsilon\to 0$ (viz. (9)) give:
$\int_{\Omega}F^{\alpha\beta}(\varepsilon)\left(e_{\alpha\|\beta}(\varepsilon;\bm{v}(\varepsilon;\bm{\eta}))-e_{\alpha\|\beta}(\varepsilon)\right)\sqrt{g(\varepsilon)}\,\mathrm{d}x$ $\to\int_{\Omega}F^{\alpha\beta}(\gamma_{\alpha\beta}(\bm{\eta})-e_{\alpha\|\beta})\sqrt{a}\,\mathrm{d}x,$
$-2\int_{\Omega}F^{\alpha 3}(\varepsilon)\left(e_{\alpha\|3}(\varepsilon;\bm{v}(\varepsilon;\bm{\eta}))-e_{\alpha\|3}(\varepsilon)\right)\sqrt{g(\varepsilon)}\,\mathrm{d}x$ $\to\int_{\Omega}F^{\alpha 3}\left(-(b_{\alpha}^{\sigma}\eta_{\sigma}+\partial_{\alpha}\eta_{3})+\dfrac{a_{\alpha\beta}}{\mu}F^{\beta 3}\right)\sqrt{a}\,\mathrm{d}x,$
$-\int_{\Omega}F^{33}(\varepsilon)\left(e_{3\|3}(\varepsilon;\bm{v}(\varepsilon;\bm{\eta}))-e_{3\|3}(\varepsilon)\right)\sqrt{g(\varepsilon)}\,\mathrm{d}x$ $\to\int_{\Omega}F^{33}\left(-\dfrac{\lambda}{\lambda+2\mu}a^{\alpha\beta}e_{\alpha\|\beta}+\dfrac{F^{33}}{\lambda+2\mu}\right)\sqrt{a}\,\mathrm{d}x,$
$-\int_{\Omega}A^{ijk\ell}(\varepsilon)e_{k\|\ell}e_{i\|j}\sqrt{g(\varepsilon)}\,\mathrm{d}x$ $\to-\int_{\omega}a^{\alpha\beta\sigma\tau}\overline{e_{\sigma\|\tau}}\,\overline{e_{\alpha\|\beta}}\sqrt{a}\,\mathrm{d}y$
$\qquad-\int_{\Omega}\left\{\dfrac{a_{\alpha\beta}}{\mu}F^{\alpha 3}F^{\beta 3}+\dfrac{|F^{33}|^{2}}{\lambda+2\mu}\right\}\sqrt{a}\,\mathrm{d}y,$
$\int_{\Omega}A^{\alpha\beta\sigma\tau}(\varepsilon)e_{\sigma\|\tau}(\varepsilon)e_{\alpha\|\beta}(\varepsilon;\bm{v}(\varepsilon;\bm{\eta}))\sqrt{g(\varepsilon)}\,\mathrm{d}x$ $\to\int_{\Omega}\{\lambda a^{\alpha\beta}a^{\sigma\tau}+\mu(a^{\alpha\sigma}a^{\beta\tau}+a^{\alpha\tau}a^{\beta\sigma})\}e_{\sigma\|\tau}\gamma_{\alpha\beta}(\bm{\eta})\sqrt{a}\,\mathrm{d}x,$
$\int_{\Omega}A^{\sigma\tau 33}(\varepsilon)e_{\sigma\|\tau}(\varepsilon)e_{3\|3}(\varepsilon;\bm{v}(\varepsilon;\bm{\eta}))\sqrt{g(\varepsilon)}\,\mathrm{d}x$ $=0,$
$4\int_{\Omega}A^{\alpha 3\sigma 3}(\varepsilon)e_{\sigma\|3}(\varepsilon)e_{\alpha\|3}(\varepsilon;\bm{v}(\varepsilon;\bm{\eta}))\sqrt{g(\varepsilon)}\,\mathrm{d}x$ $\to\int_{\Omega}F^{\alpha 3}(b_{\alpha}^{\sigma}\eta_{\sigma}+\partial_{\alpha}\eta_{3})\sqrt{a}\,\mathrm{d}x,$
$\int_{\Omega}A^{33\sigma\tau}(\varepsilon)e_{3\|3}(\varepsilon)e_{\sigma\|\tau}(\varepsilon;\bm{v}(\varepsilon;\bm{\eta}))\sqrt{g(\varepsilon)}\,\mathrm{d}x$ $\to\int_{\Omega}\lambda a^{\sigma\tau}\left(-\dfrac{\lambda}{\lambda+2\mu}a^{\alpha\beta}e_{\alpha\|\beta}+\dfrac{F^{33}}{\lambda+2\mu}\right)\gamma_{\sigma\tau}(\bm{\eta})\sqrt{a}\,\mathrm{d}x,$
$\int_{\Omega}A^{3333}(\varepsilon)e_{3\|3}(\varepsilon)e_{3\|3}(\varepsilon;\bm{v}(\varepsilon;\bm{\eta}))\sqrt{g(\varepsilon)}\,\mathrm{d}x$ $=0.$
Therefore, combining (10) with the asymptotics above and passing to the averages where possible gives:
$\int_{\omega}a^{\alpha\beta\sigma\tau}\overline{e_{\sigma\|\tau}}\left(\gamma_{\alpha\beta}(\bm{\eta})-\overline{e_{\alpha\|\beta}}\right)\sqrt{a}\,\mathrm{d}y\geq\int_{\omega}\left(\int_{-1}^{1}F^{\alpha\beta}-\dfrac{\lambda}{\lambda+2\mu}a^{\alpha\beta}F^{33}\,\mathrm{d}x_{3}\right)\left(\gamma_{\alpha\beta}(\bm{\eta})-\overline{e_{\alpha\|\beta}}\right)\sqrt{a}\,\mathrm{d}y,$
which is the sought inequality.
(vii) The subsequence $\{\bm{u}(\varepsilon)\}_{\varepsilon>0}$ found in part (ii) is such that:
$\varepsilon\bm{u}(\varepsilon)$ $\rightharpoonup 0,\quad\textup{ in }\bm{H}^{1}(\Omega),$
$\partial_{3}u_{\alpha}(\varepsilon)$ $\rightharpoonup 0,\quad\textup{ in }L^{2}(\Omega),$
as $\varepsilon\to 0$. Furthermore, the weak limits $e_{i\|j}$ are independent of the transverse variable $x_{3}$.
The proof is identical to the proof of part (vi) of Theorem 5.6-1 of [8] and is, for this reason, omitted.
(viii) The following convergences, that are so far only known to be weak, are actually strong:
$e_{i\|j}(\varepsilon)$ $\to e_{i\|j},\quad\textup{ in }L^{2}(\Omega),$
$\varepsilon\bm{u}(\varepsilon)$ $\to\bm{0},\quad\textup{ in }\bm{H}^{1}(\Omega),$
$\gamma_{\alpha\beta}\left(\overline{\bm{u}(\varepsilon)}\right)$ $\to\overline{e_{\alpha\|\beta}},\quad\textup{ in }L^{2}(\Omega),$
$\overline{\bm{u}(\varepsilon)}$ $\to\bm{\zeta},\quad\textup{ in }\bm{U}_{M}^{\sharp}(\Omega).$
Consider the restriction of the linearised strain tensor to the space $\bm{V}(\omega)$, that we still denote and define by:
$\gamma_{\alpha\beta}:\bm{V}(\omega)\to L^{2}(\omega).$
Combining the linearity and continuity of one such operator with the fact that the assumption that the semi-norm $|\cdot|_{\omega}^{M}$ is actually a norm over $\bm{V}(\omega)$ (since the generalised membrane shell under consideration is of the first kind), allows us to apply Theorem 3.1-1 in [10], so as to infer that there exists a unique linear and continuous extension that we denote and define by:
$\gamma_{\alpha\beta}^{\sharp}:\bm{V}_{M}^{\sharp}(\omega)\to L^{2}(\omega).$
Let us recall that the following convergences were established in part (i) and (v), respectively:
[…]
The linearity and continuity of $\gamma_{\alpha\beta}^{\sharp}$ together with the uniqueness of the weak limit (cf., e.g., [2]), allow us to infer that:
$\gamma_{\alpha\beta}^{\sharp}(\bm{\zeta})=\overline{e_{\alpha\|\beta}}\textup{in }L^{2}(\omega).$
In light of the definition of $\bm{U}_{M}^{\sharp}(\omega)$, we have that in correspondence of the vector field $\bm{\zeta}\in\bm{U}_{M}^{\sharp}(\omega)$ satisfying the variational inequalities in part (vi), it is possible to individuate a sequence $\{\bm{\zeta}_{\ell}\}_{\ell\geq 1}\subset\bm{U}(\omega)$ such that:
$\bm{\zeta}_{\ell}\to\bm{\zeta},\quad\textup{ in }\bm{U}_{M}^{\sharp}(\omega)\textup{ as }\ell\to\infty,$
so as to obtain at once that:
(11) $\gamma_{\alpha\beta}^{\sharp}(\bm{\zeta}_{\ell})\to\overline{e_{\alpha\|\beta}},\quad\textup{ in }L^{2}(\omega)\textup{ as }\ell\to\infty.$
The same computations as in part (vi) and the properties of the extension $\gamma_{\alpha\beta}^{\sharp}$ observed beforehand lead to obtain the following estimates
(12) 00 0 $\leq\dfrac{\sqrt{g_{0}}}{C_{e}c_{0}^{2}}\limsup_{\varepsilon\to 0}\left(\sum_{i,j}\|e_{i\|j}(\varepsilon)-e_{i\|j}\|_{L^{2}(\Omega)}^{2}\right)\leq\int_{\omega}a^{\alpha\beta\sigma\tau}\overline{e_{\sigma\|\tau}}\left(\gamma_{\alpha\beta}^{\sharp}(\bm{\zeta}_{\ell})-\overline{e_{\alpha\|\beta}}\right)\sqrt{a}\,\mathrm{d}y$
$\quad-\int_{\omega}\left(\int_{-1}^{1}F^{\alpha\beta}-\dfrac{\lambda}{\lambda+2\mu}a^{\alpha\beta}F^{33}\,\mathrm{d}x_{3}\right)\left(\gamma_{\alpha\beta}^{\sharp}(\bm{\zeta}_{\ell})-\overline{e_{\alpha\|\beta}}\right)\sqrt{a}\,\mathrm{d}y,$
which are valid for each $\ell\geq 1$. Thanks to (11), we obtain that the right-hand side in (12) tends to zero as $\ell\to\infty$. We have thus established that:
$\limsup_{\varepsilon\to 0}\|e_{i\|j}(\varepsilon)-e_{i\|j}\|_{L^{2}(\Omega)}=0,\quad\textup{ for all }i,j.$
The remaining strong convergences can be established in the same fashion of part (vii) in Theorem 5.6-1 of [8]. In particular, we have obtained that the elements $\overline{\bm{u}(\varepsilon)}\in\bm{U}(\omega)$ strongly converge to $\bm{\zeta}\in\bm{U}_{M}^{\sharp}(\omega)$ as $\varepsilon\to 0$.
(ix) The limit $\bm{\zeta}\in\bm{U}_{M}^{\sharp}(\omega)$ found in part (viii) is the unique solution of Problem $\mathcal{P}_{M}^{\sharp}(\omega)$. Consequently, the entire sequence $\{\overline{\bm{u}(\varepsilon)}\}_{\varepsilon>0}$ converges to $\bm{\zeta}$ in $\bm{V}_{M}^{\sharp}(\omega)$ as $\varepsilon\to 0$.
Fix $\bm{\eta}\in\bm{U}(\omega)$, and recall that $\overline{\bm{u}(\varepsilon)}\in\bm{U}(\omega)$. The strong convergences obtained upon completion of part (viii) give that:
(13) $\int_{\omega}a^{\alpha\beta\sigma\tau}\gamma_{\sigma\tau}\left(\overline{\bm{u}(\varepsilon)}\right)\gamma_{\alpha\beta}(\bm{\eta})\sqrt{a}\,\mathrm{d}y$ $\to\int_{\omega}a^{\alpha\beta\sigma\tau}\overline{e_{\sigma\|\tau}}\gamma_{\alpha\beta}(\bm{\eta})\sqrt{a}\,\mathrm{d}y,$
$\int_{\omega}a^{\alpha\beta\sigma\tau}\gamma_{\sigma\tau}\left(\overline{\bm{u}(\varepsilon)}\right)\gamma_{\alpha\beta}\left(\overline{\bm{u}(\varepsilon)}\right)\sqrt{a}\,\mathrm{d}y$ $\to\int_{\omega}a^{\alpha\beta\sigma\tau}\overline{e_{\sigma\|\tau}}\,\overline{e_{\alpha\|\beta}}\sqrt{a}\,\mathrm{d}y.$
The convergences in (13) and the conclusion of part (vi) imply, on the one hand, that:
(14) $\lim_{\varepsilon\to 0}$ $B_{M}\left(\overline{\bm{u}(\varepsilon)},\bm{\eta}-\overline{\bm{u}(\varepsilon)}\right)=\int_{\omega}a^{\alpha\beta\sigma\tau}\overline{e_{\sigma\|\tau}}\left(\gamma_{\alpha\beta}(\bm{\eta})-\overline{e_{\alpha\|\beta}}\right)\sqrt{a}\,\mathrm{d}y$
$\geq\int_{\omega}\varphi^{\alpha\beta}\left(\gamma_{\alpha\beta}(\bm{\eta})-\overline{e_{\alpha\|\beta}}\right)\sqrt{a}\,\mathrm{d}y=\lim_{\varepsilon\to 0}L_{M}\left(\bm{\eta}-\overline{\bm{u}(\varepsilon)}\right)=L_{M}^{\sharp}(\bm{\eta}-\bm{\zeta}).$
On the other hand, it is also true that:
(15) $\lim_{\varepsilon\to 0}B_{M}\left(\overline{\bm{u}(\varepsilon)},\bm{\eta}-\overline{\bm{u}(\varepsilon)}\right)=B_{M}^{\sharp}(\bm{\zeta},\bm{\eta}-\bm{\zeta}).$
Therefore, combining (14) and (15), we have that the definition of continuous extensions $B_{M}^{\sharp}$ and $L_{M}^{\sharp}$ gives:
$B_{M}^{\sharp}(\bm{\zeta},\bm{\eta}-\bm{\zeta})\geq L_{M}^{\sharp}(\bm{\eta}-\bm{\zeta}),\quad\textup{ for all }\bm{\eta}\in\bm{U}_{M}^{\sharp}(\omega).$
Noting that the set $\bm{U}_{M}^{\sharp}(\omega)$ is non-empty, closed and convex (viz. part (ii)), noting that a straightforward density argument gives that the bilinear form $B_{M}^{\sharp}$ and $L_{M}^{\sharp}$ satisfy the assumptions of Theorem 6.1-2 in [10], we infer that there exists a unique solution for Problem $\mathcal{P}_{M}^{\sharp}(\omega)$, and that the entire sequence $\{\overline{\bm{u}(\varepsilon)}\}_{\varepsilon>0}$ converges to $\bm{\zeta}$ in $\bm{V}_{M}^{\sharp}(\omega)$ as $\varepsilon\to 0$.
(x) Conclusion of the proof.
The proof of parts (ix), (x) and (xi) of Theorem 5.6-1 in [8] do not need any adaptation and their conclusions thus follow, completing the proof. ∎
6. Justification of Koiter’s model for linearly elastic generalised membrane shells of the first kind
Upon completing the proof of Theorem 5.1, it remains to show that the solution $\bm{u}^{\varepsilon}$ of Problem $\mathcal{P}(\Omega^{\varepsilon})$ asymptotically behaves like the solution of Koiter’s model. We recall that the contrivance of Koiter’s model, which is due to Koiter [29, 30], is based on the Kirchhoff-Love assumptions, which are assumptions of geometrical nature, and on the assumption that the estimates derived by F. John [25, 26] are valid. The most remarkable feature of Koiter’s model is that it is a two-dimensional model that is meant to approximate the original three-dimensional linearly elastic energy regardless the type of shell under consideration. For a more detailed description of Koiter’s model we refer the reader to Chapter 7 of [8].
The functional space where Koiter’s model is defined is given by:
$\bm{V}_{K}(\omega):=\{\bm{\eta}=(\eta_{i})\in H^{1}(\omega)\times H^{1}(\omega)\times H^{2}(\omega);\eta_{i}=\partial_{\nu}\eta_{3}=0\textup{ on }\gamma_{0}\}.$
For each $\varepsilon>0$, the free energy associated with Koiter’s model for a linearly elastic generalised membrane shell of the first kind takes the following form:
$J_{K}^{\varepsilon}(\bm{\eta}):=\dfrac{\varepsilon}{2}\int_{\omega}a^{\alpha\beta\sigma\tau}\gamma_{\sigma\tau}(\bm{\eta})\gamma_{\alpha\beta}(\bm{\eta})\sqrt{a}\,\mathrm{d}y+\dfrac{\varepsilon^{3}}{6}\int_{\omega}a^{\alpha\beta\sigma\tau}\rho_{\sigma\tau}(\bm{\eta})\rho_{\alpha\beta}(\bm{\eta})\sqrt{a}\,\mathrm{d}y-\varepsilon\int_{\omega}\varphi^{\alpha\beta}\gamma_{\alpha\beta}(\bm{\eta})\sqrt{a}\,\mathrm{d}y,$
for all $\bm{\eta}=(\eta_{i})\in\bm{V}_{K}(\omega)$.
In the case where the shell is required to remain confined in a prescribed half-space, the competitors to the role of minimiser for the energy functional $J_{K}^{\varepsilon}$ are to be sought in the following non-empty, closed and convex subset of the space $\bm{V}_{K}(\omega)$ introduced beforehand:
$\bm{U}_{K}(\omega):=\{\bm{\eta}=(\eta_{i})\in\bm{V}_{K}(\omega);(\bm{\theta}+\eta_{i}\bm{a}^{i})\cdot\bm{q}\geq 0\textup{ a.e. in }\omega\}.$
Let us now recall the statement of Theorem 4.1 in [16] concerning the identification of a two-dimensional limit model for linearly elastic generalised membrane shells of the first kind subjected to remaining confined in a half-space whose orthogonal complement is the given unit-vector $\bm{q}$.
Theorem 6.1.
Let $\omega$ be a domain in $\mathbb{R}^{2}$ and let $\bm{\theta}\in\mathcal{C}^{3}(\overline{\omega};\mathbb{E}^{3})$ be an injective immersion such that $\min_{y\in\overline{\omega}}(\bm{\theta}\cdot\bm{q})=d>0$. Consider a family of generalised membrane shells of the first kind with thickness $2\varepsilon$ approaching zero and with each having the same middle surface $\bm{\theta}(\overline{\omega})$, and assume that each shell is subject to a boundary condition of place along a portion of its lateral face, whose middle curve is the set $\bm{\theta}(\gamma_{0})$.
Define the spaces
$\bm{V}(\omega)$ :=$:=\{\bm{\eta}=(\eta_{i})\in\bm{H}^{1}(\omega);\bm{\eta}=\bm{0}\textup{ on }\gamma_{0}\},$
$\bm{V}_{M}^{\sharp}(\omega)$ $:=\textup{completion of $\bm{V}(\omega)$ with respect to $\left|\cdot\right|_{\omega}^{M}$}.$
For each $\varepsilon>0$, let $\bm{\zeta}_{K}^{\varepsilon}$ denote the unique minimiser of the energy functional $J_{K}^{\varepsilon}$ over the set $\bm{U}_{K}(\omega)$.
Then the following convergence holds:
$\bm{\zeta}_{K}^{\varepsilon}\to\bm{\zeta},\quad\textup{ in }\bm{V}_{M}^{\sharp}(\omega)\textup{ as }\varepsilon\to 0,$
where $\bm{\zeta}$ denote the unique solution of the following variational problem: Find
$\bm{\zeta}\in\bm{U}^{\sharp}(\omega):=\textup{closure of $\bm{U}_{K}(\omega)$ with respect to $\left|\cdot\right|_{\omega}^{M}$}$
that satisfies the following variational inequalities:
$B_{M}^{\sharp}(\bm{\zeta},\bm{\eta}-\bm{\zeta})\geq L_{M}^{\sharp}(\bm{\eta}-\bm{\zeta})\quad\textup{ for all }\bm{\eta}=(\eta_{i})\in\bm{U}^{\sharp}(\omega),$
where $B_{M}^{\sharp}(\cdot,\cdot)$ and $L_{M}^{\sharp}$ designate the unique continuous linear extensions from $\bm{V}(\omega)$ to $\bm{V}_{M}^{\sharp}(\omega)$ of the bilinear form $B_{M}(\cdot,\cdot)$, and of the linear form $L_{M}$ defined by
$L_{M}(\bm{\eta}):=\int_{\omega}\varphi^{\alpha\beta}\gamma_{\alpha\beta}(\bm{\eta})\sqrt{a}\,\mathrm{d}y\quad\textup{ for all }\bm{\eta}=(\eta_{i})\in\bm{V}(\omega).$
In order to fully justify Koiter’s model for linearly elastic generalised membrane shells of the first kind subjected to remaining confined in a prescribed half-space, we are left to show that:
(16) $\bm{U}_{M}^{\sharp}(\omega):=\overline{\bm{U}(\omega)}^{|\cdot|_{\omega}^{M}}=\bm{U}^{\sharp}(\omega):=\overline{\bm{U}_{K}(\omega)}^{|\cdot|_{\omega}^{M}}.$
To this aim, we will need to resort to the following density result, established by P. Doktor [21] in a more general form (see also [22]).
Theorem 6.2.
Let $V:=\{\eta\in H^{1}(\omega);\eta=0\textup{ on }\gamma_{0}\}$. Then, the space $\mathcal{C}^{\infty}(\overline{\omega})\cap V$ is dense in $V$ with respect to the norm $\|\cdot\|_{H^{1}(\omega)}$. ∎
The next theorem, that constitutes the second main result of this paper, is devoted to establishing the set equality (16) under the assumption that the domain $\omega$ is a rectangle. Let $\hat{\omega}$ denote an extension of $\omega$ obtained by first reflecting a strip of given width with respect to a pair of parallel edges and by then reflecting the strip of the same width associated with the obtained intermediary extension with respect to the remaining pair of parallel edges. In this case, we can extend the injective immersion $\bm{\theta}$ to a new injective mapping $\hat{\bm{\theta}}:\overline{\hat{\omega}}\to\mathbb{E}^{3}$ as follows: First, we reflect $\bm{\theta}$ along the edges of the rectangle that belong to $\gamma\setminus\gamma_{0}$, and then we smoothly extend $\bm{\theta}$ via the Whitney extension theorem (cf., e.g., Theorem 2.3.6 of [24] and Lemma 6.3 in [36] for a similar argument) along the edges that belong to $\gamma_{0}$. In light of the latter considerations, we have that $\hat{\bm{\theta}}\in\mathcal{C}^{3}(\overline{\hat{\omega}}\setminus\gamma;\mathbb{E}^{3})\cap\mathcal{C}^{0}(\overline{\hat{\omega}};\mathbb{E}^{3})$.
Likewise, we extend the contravariant basis $\{\bm{a}^{i}\}_{i=1}^{3}$ associated with the injective immersion $\bm{\theta}$ by reflection along the edges of the boundary $\gamma\setminus\gamma_{0}$ and by smooth extension along the edges of the rectangle that belong to $\gamma_{0}$. We denote the extension of the vector field $\bm{a}^{i}$ by $\hat{\bm{a}}^{i}$ and we choose the unit-vector $\bm{q}$ in a way that
(17) $\hat{\bm{a}}^{i}(\hat{y})\cdot\bm{q}=\bm{a}^{i}(y)\cdot\bm{q},$
for all $y\in\omega$ and $\hat{y}\in\hat{\omega}\setminus\omega$ that is antipodal to $y$ with respect to some point $y_{0}\in\gamma\setminus\gamma_{0}$, and for each $1\leq i\leq 3$. A geometrical interpretation for the condition (17) is illustrated in Figure 1 below.
Figure 1. An example where the condition (17) holds. The figure here represented is a cross section of a portion of a cylinder, whose extension by reflection along one of the straight edges of its boundary is denoted by the dotted pattern. The unit-vector $\bm{q}$ identifying the orthogonal complement to the plane identifying the half-space where the shell has to remain confined is given and it is equal to $(0,0,1)$. The vector $\bm{a}^{3}(y)$, denoted by a solid line, is reflected onto the corresponding vector $\hat{\bm{a}}^{3}(\hat{y})$, denoted by a dotted line, where $\hat{y}\in\overline{\hat{\omega}}\setminus\omega$ is the unique antipodal point with respect to the straight edge of the boundary $\gamma$ that corresponds to an appropriate point $y\in\omega$. We observe that $\bm{a}^{3}(y)\cdot\bm{q}=\hat{\bm{a}}^{3}(\hat{y})\cdot\bm{q}$, even though $\bm{a}^{3}(y)\neq\hat{\bm{a}}^{3}(\hat{y})$. Other examples of parametrisation of surfaces $\bm{\theta}$ and unit-vectors $\bm{q}$ for which this extension machinery is possible are described in Figure 5. The remaining vectors of the contravariant basis under consideration can be extended likewise.
Note that, in general, we have that $\bm{a}^{i}(y)\neq\hat{\bm{a}}^{i}(\hat{y})$, although the mappings
$y\in\overline{\omega}\mapsto\hat{\bm{a}}^{i}(y)\cdot\bm{q},\quad 1\leq i\leq 3,$
are continuous. These observations lead to the formulation of the following density theorem, which constitutes the second new result in this paper and that is critical for establishing the justification of Koiter’s model for linearly elastic generalised membrane shells subjected to remaining confined in a half-space.
Theorem 6.3.
Let $\bm{\theta}\in\mathcal{C}^{3}(\overline{\omega};\mathbb{E}^{3})$ be an injective immersion such that $\min_{y\in\overline{\omega}}(\bm{\theta}\cdot\bm{q})=d>0$.
Assume that $\omega$ satisfies one and only one of the following:
  • $(a)$
    $\gamma_{0}=\gamma$;
  • $(b)$
    $\omega$ is a rectangle and $\gamma_{0}$ is equal to the union of edges of $\gamma$.
Assume that the unit-vector $\bm{q}\in\mathbb{E}^{3}$ is such that (17) holds, i.e.,
$\hat{\bm{a}}^{i}(\hat{y})\cdot\bm{q}=\bm{a}^{i}(y)\cdot\bm{q},$
for all $y\in\omega$ and $\hat{y}\in\hat{\omega}\setminus\omega$ that is antipodal to $y$ with respect to some point $y_{0}\in\gamma\setminus\gamma_{0}$, and for each $1\leq i\leq 3$. Then, we have that:
$\bm{U}_{M}^{\sharp}(\omega)=\bm{U}^{\sharp}(\omega).$
Proof.
The sought set equality will be shown by resorting to a density argument, aiming to construct smaller and smaller dense subsets of $\bm{U}_{M}(\omega)$ with respect to the norm $\|\cdot\|_{\bm{H}^{1}(\omega)}$. The fact that the linearly elastic generalised membrane shell under consideration is of the first kind renders the semi-norm $|\cdot|_{\omega}^{M}$ a norm over the space $\bm{V}(\omega)$ and, moreover, we have that there exists a constant $C=C(\bm{\theta},\omega,\lambda,\mu)>0$ such that:
(18) $|\bm{\eta}|_{\omega}^{M}\leq C\|\bm{\eta}\|_{\bm{H}^{1}(\omega)},\quad\textup{for all }\bm{\eta}\in\bm{V}(\omega).$
The proof is broken into four parts numbered (i)–(iv).
(i) The set $\bm{U}_{M}^{(1)}(\omega):=\bm{U}(\omega)\cap\bm{L}^{\infty}(\omega)$ is dense in the set $\bm{U}(\omega)$ with respect to the norm $\|\cdot\|_{\bm{H}^{1}(\omega)}$.
Let $\bm{\eta}\in\bm{U}(\omega)$ be given. For each integer $k\geq 1$, define the vector field $\bm{\eta}^{(k)}=\eta_{i}^{(k)}$ in a way such that its components satisfy the following relations:
$\eta_{i}^{(k)}(y)=\begin{cases}\eta_{i}(y)&,\textup{ for a.a. }y\in\omega\textup{ such that }|\eta_{j}(y)\bm{a}^{j}(y)|\leq k,\\ \\ \dfrac{k\eta_{i}(y)}{|\eta_{j}(y)\bm{a}^{j}(y)|}&,\textup{ for a.a. }y\in\omega\textup{ such that }|\eta_{j}(y)\bm{a}^{j}(y)|>k.\end{cases}$
Therefore, we have that $|\eta_{i}^{(k)}\bm{a}^{i}|\leq k$ if $|\eta_{j}\bm{a}^{j}|\leq k$, and that $|\eta_{i}^{(k)}\bm{a}^{i}|\equiv k$ if $|\eta_{j}\bm{a}^{j}|>k$.
An application of the Cahuchy-Schwarz inequality gives
$|\eta_{i}^{(k)}|\leq|\eta_{j}^{(k)}\bm{a}^{j}||\bm{a}_{i}|\leq k\max_{1\leq\ell\leq 3}\|\bm{a}_{\ell}\|_{\bm{\mathcal{C}}^{1}(\overline{\omega})},\quad\textup{ for a.a. }y\in\omega,$
thus showing that $\eta_{i}^{(k)}\in L^{\infty}(\omega)$. If $y\in\omega$ is such that $|\eta_{j}(y)\bm{a}^{j}(y)|\leq k$, then
$(\bm{\theta}(y)+\eta_{i}^{(k)}(y)\bm{a}^{i}(y))\cdot\bm{q}=(\bm{\theta}(y)+\eta_{i}(y)\bm{a}^{i}(y))\cdot\bm{q}\geq 0,$
whereas, if $|\eta_{j}(y)\bm{a}^{j}(y)|>k$, then
$(\bm{\theta}(y)+\eta_{i}^{(k)}(y)\bm{a}^{i}(y))\cdot\bm{q}=\left(\bm{\theta}(y)+\dfrac{k}{|\eta_{j}(y)\bm{a}^{j}(y)|}\eta_{i}(y)\bm{a}^{i}(y)\right)\cdot\bm{q}\geq\left(1-\dfrac{k}{|\eta_{j}(y)\bm{a}^{j}(y)|}\right)\bm{\theta}\cdot\bm{q}>0.$
Observe that we can write
$\eta_{i}^{(k)}=g^{(k)}(u)\eta_{i},$
where $u:=|\eta_{j}\bm{a}^{j}|^{2}=\eta_{j}\eta_{\ell}a^{j\ell}$ and the function $g^{(k)}:[0,\infty)\to\mathbb{R}$ is defined by:
$g^{(k)}(u):=\begin{cases}1&,\textup{ if }0\leq u\leq k^{2},\\ \\ \dfrac{k}{\sqrt{u}}&,\textup{ if }u>k^{2}.\end{cases}$
The function $g^{(k)}$ is continuous and it is differentiable in the classical sense in $(0,k^{2})\cup(k^{2},\infty)$, with derivative given by:
$\dfrac{\,\mathrm{d}g^{(k)}}{\,\mathrm{d}u}(u)=\begin{cases}0&,\textup{ if }0<u<k^{2},\\ \\ \dfrac{-k}{2u^{3/2}}&,\textup{ if }u>k^{2}.\end{cases}$
Since $\eta_{j}\in H^{1}(\omega)$, then the Sobolev embedding theorem in the limiting case (Corollary 9.11 in [2]) ensures that $\eta_{j}\in L^{r}(\omega)$ for all $2\leq r<\infty$. An application of the generalised Hölder’s inequality inequality (cf., e.g., Exercise 3.7 in [6]) thus gives that $u\in W^{1,q}(\omega)$, for some $21\leq q<$ 2. Since
$\left|\frac{\,\mathrm{d}g^{(k)}}{\,\mathrm{d}u}\right|<\frac{1}{2k^{2}},$
we have obtained that $g^{(k)}\in W^{1,\infty}(0,\infty)$ and, therefore, that $g^{(k)}$ is Lipschitz continuous (cf., e.g., Proposition 8.4 of [2]). Therefore, we are in a position to apply Theorem 2.2 in [35], so as to obtain that:
$g^{(k)}(u)\in W^{1,q}(\omega),\quad 1<q<2.$
Furthermore, the weak derivatives of $g^{(k)}(u)$ are given by the standard chain rule (cf., Theorem 2.1 in [35]):
$\partial_{\beta}(g^{(k)}(u))=\begin{cases}0&,\textup{ if }0<u<k^{2},\\ \\ -\dfrac{k}{2u^{3/2}}\partial_{\beta}u&,\textup{ if }u>k^{2}.\end{cases}$
In order to show that $\eta_{i}^{(k)}\in H^{1}(\omega)$, we consider a subset $\tilde{\omega}\subset\subset\omega$, we let $h\in\mathbb{R}^{2}$ be such that $|h|<\textup{dist}(\tilde{\omega},\gamma)$. An application of the Taylor-MacLaurin formula (cf., e.g., Theorem 7.9-1(c) in [10]) gives:
$\left(\int_{\tilde{\omega}}\left|\eta_{i}^{(k)}(y+h)-\eta_{i}^{(k)}(y)\right|^{2}\,\mathrm{d}y\right)^{1/2}$
$\leq\left(\int_{\tilde{\omega}}\left|g^{(k)}(u)(y+h)-g^{(k)}(u)(y)\right|^{2}|\eta_{i}(y+h)|^{2}\,\mathrm{d}y\right)^{1/2}$
$\qquad+\left(\int_{\tilde{\omega}}\left|g^{(k)}(u)(y)\right|^{2}|\eta_{i}(y+h)-\eta_{i}(y)|^{2}\,\mathrm{d}y\right)^{1/2}$
$\leq\left(\dfrac{1}{2k^{2}}\|\eta_{i}\|_{L^{2}(\omega)}+\|\nabla\eta_{i}\|_{\bm{L}^{2}(\omega)}\right)|h|.$
We are thus in a position to apply Proposition 9.3 in [2] to infer that $\eta_{i}^{(k)}\in H^{1}(\omega)$. In order to compute the weak derivatives of $\eta_{i}^{(k)}$, we proceed by density. In light of Theorem 6.2, we can find a sequence $\{\eta_{i,\ell}\}_{\ell\geq 1}\subset\mathcal{C}^{\infty}(\overline{\omega})$ such that $\eta_{i,\ell}=0$ on $\gamma_{0}$, and such that
(19) $\eta_{i,\ell}\to\eta_{i},\quad\textup{ in }H^{1}(\omega)\textup{ as }\ell\to\infty.$
Evaluate
$\|g^{(k)}(u)\eta_{j,\ell}-g^{(k)}(u)\eta_{j}\|_{L^{2}(\omega)}\leq\|g^{(k)}(u)\|_{L^{\infty}(\omega)}\|\eta_{j,\ell}-\eta_{j}\|_{L^{2}(\omega)},$
and note that, thanks to (19), the right-hand side approaches zero as $\ell\to\infty$. In conclusion, we have obtained that:
(20) $g^{(k)}(u)\eta_{j,\ell}\to g^{(k)}(u)\eta_{j},\quad\textup{ in }L^{2}(\omega)\textup{ as }\ell\to\infty.$
Let us now estimate the term:
(21) $\|g^{(k)}(u)\partial_{\beta}\eta_{i,\ell}-g^{(k)}(u)\partial_{\beta}\eta_{i}\|_{L^{2}(\omega)}\leq\|g^{(k)}(u)\|_{L^{\infty}(\omega)}\|\partial_{\beta}\eta_{i,\ell}-\partial_{\beta}\eta_{i}\|_{L^{2}(\omega)},$
and note that the right-hand side approaches zero as $\ell\to\infty$ thanks to (19) and the generalised Hölder’s inequality recalled beforehand. In conclusion, we have obtained that:
(22) $g^{(k)}(u_{\ell})\partial_{\beta}\eta_{i,\ell}\to g^{(k)}(u)\partial_{\beta}\eta_{i},\quad\textup{ in }L^{2}(\omega)\textup{ as }\ell\to\infty.$
Recall that $\partial_{\beta}u\in L^{q}(\omega)$ for some $21<q<$ 2, define the positive number $r:=3q/(2q-2)$, define the positive number $p:=3q/(2q+1)$ and note that $21<p<$ 2. Besides, we have that:
$\dfrac{1}{p}=\dfrac{1}{q}+\dfrac{1}{r}.$
Let us now estimate the term:
(23) $\|\partial_{\beta}g^{(k)}(u)\eta_{i,\ell}-\partial_{\beta}g^{(k)}(u)\eta_{i}\|_{L^{p}(\omega)}\leq\|\partial_{\beta}g^{(k)}(u)\|_{L^{q}(\omega)}\|\eta_{i,\ell}-\eta_{i}\|_{L^{r}(\omega)}\leq\dfrac{\|\partial_{\beta}u\|_{L^{q}(\omega)}}{2k^{2}}\|\eta_{i,\ell}-\eta_{i}\|_{L^{r}(\omega)},$
and note that the right-hand side approaches zero as $\ell\to\infty$ thanks to (19). In conclusion, we have obtained that:
(24) $\partial_{\beta}g^{(k)}(u)\eta_{i,\ell}\to\partial_{\beta}g^{(k)}(u)\eta_{i},\quad\textup{ in }L^{p}(\omega)\textup{ as }\ell\to\infty.$
In light of (20)–(24), we have obtained that
(25) $g^{(k)}(u)\eta_{j,\ell}\to g^{(k)}(u)\eta_{j},\quad\textup{ in }W^{1,p}(\omega)\textup{ as }\ell\to\infty.$
Thanks to (20)–(25), we have that $\eta_{i}^{(k)}\in W^{1,p}(\omega)$ with $21<p<$ 2. Besides, we are able to express the weak derivatives $\partial_{\beta}\eta_{i}^{(k)}$ by means of the classical chain rule:
$\partial_{\beta}\eta_{i}^{(k)}=\partial_{\beta}(g^{(k)}(u))\eta_{i}+g^{(k)}(u)\partial_{\beta}\eta_{i}.$
In light of (25), we can thus assert the convergence
(26) $g^{(k)}(u)\eta_{i,\ell}\to g^{(k)}(u)\eta_{i},\quad\textup{ in }L^{p}(\gamma_{0})\textup{ as }\ell\to\infty.$
Since $g^{(k)}(u)\in W^{1,q}(\omega)$ for some $21<q<$ 2, we can find a sequence $\{\psi_{n}\}_{n\geq 0}\subset\mathcal{C}^{\infty}(\overline{\omega})$ such that $\psi_{n}\to g^{(k)}(u)$ in $W^{1,q}(\omega)$ as $n\to\infty$. Therefore, we have that:
(27) $\psi_{n}\eta_{i,\ell}\to g^{(k)}(u)\eta_{i,\ell},\quad\textup{ in }W^{1,q}(\omega)\textup{ as }n\to\infty.$
Since $\psi_{n}$ and $\eta_{i,\ell}$ are both smooth and since $\eta_{i,\ell}$ vanishes along $\gamma_{0}$, we have that $\psi_{n}\eta_{i,\ell}\equiv 0$ on $\gamma_{0}$. Therefore, combining the Hölder inequality, (26) and (27), we obtain that $\eta_{i}^{(k)}=0$ on $\gamma_{0}$.
It remains to show that $\|\bm{\eta}^{(k)}-\bm{\eta}\|_{\bm{H}^{1}(\omega)}\to 0$ as $k\to\infty$. To begin with, we observe that the relations
[…]
imply that
$\|\eta_{i}^{(k)}-\eta_{i}\|_{L^{2}(\omega)}\leq\|\eta_{i}\|_{L^{2}(\{|\eta_{j}\bm{a}^{j}|>k\})},$
so that the following convergence holds:
$\|\eta_{i}^{(k)}-\eta_{i}\|_{L^{2}(\omega)}\to 0,\quad\textup{ as }k\to\infty.$
Second, we observe that the definition of the functions $g^{(k)}$, the expressions of the weak derivatives $\partial_{\beta}\eta_{i}^{(k)}$, and the estimates established above, altogether show that
(28) $\partial_{\beta}\eta_{i}^{(k)}-\partial_{\beta}\eta_{i}=(g^{(k)}(u)-1)\partial_{\beta}\eta_{i}+\partial_{\beta}g^{(k)}(u)\eta_{i}.$
Since we have that
$\|(g^{k}(u)-1)\partial_{\beta}\eta_{\alpha}\|_{L^{2}(\omega)}$ $=\left\|\left(\frac{k}{|\eta_{j}\bm{a}^{j}|}-1\right)\partial_{\beta}\eta_{\alpha}\right\|_{L^{2}(\{|\eta_{j}\bm{a}^{j}|>k\})},$
[…]
$\|\partial_{\beta}g^{(k)}(u)\eta_{i}\|_{L^{2}(\omega)}$ $\leq\dfrac{\max_{1\leq i\leq 3}\left(\|\bm{a}_{i}\|_{\bm{\mathcal{C}}^{0}(\overline{\omega})}\right)}{2k}\|\partial_{\beta}a^{j\ell}\|_{\mathcal{C}^{0}(\overline{\omega})}\|\eta_{j}\eta_{\ell}\|_{L^{2}(\{|\eta_{j}\bm{a}^{j}|>k\})}$
$\quad+\max_{1\leq i\leq 3}\left(\|\bm{a}_{i}\|_{\bm{\mathcal{C}}^{0}(\overline{\omega})}\right)\|\bm{a}^{j}\|_{\bm{\mathcal{C}}^{0}(\overline{\omega})}\|\partial_{\beta}\eta_{j}\|_{L^{2}(\{|\eta_{j}\bm{a}^{j}|>k\})},$
the conclusion immediately follows.
(ii) Define the set:
$\bm{U}_{M}^{(2)}(\omega)$ :=$:=\{\bm{\eta}=(\eta_{i})\in\bm{U}_{M}^{(1)}(\omega);\textup{ there exists a number }\delta(\bm{\eta})>0$
[…]
Then, we have that the set $\bm{U}_{M}^{(2)}(\omega)$ is dense in the set $\bm{U}_{M}^{(1)}(\omega)$ with respect to the norm $\|\cdot\|_{\bm{H}^{1}(\omega)}$.
Let $\bm{\eta}\in\bm{U}_{M}^{(1)}(\omega)$ be given. For each integer $k\geq 1$, define the function $f^{(k)}:\overline{\omega}\to\mathbb{R}$ by:
$f^{(k)}(y):=\begin{cases}0&,\textup{ if }0\leq\textup{dist}(y,\gamma_{0})\leq\dfrac{1}{k},\\ \\ k\textup{dist}(y,\gamma_{0})-1&,\textup{ if }\dfrac{1}{k}\leq\textup{dist}(y,\gamma_{0})\leq\dfrac{2}{k},\\ \\ 1&,\textup{ if }0\leq\textup{dist}(y,\gamma_{0})>\dfrac{2}{k}.\end{cases}$
It is immediate to see that $f^{(k)}\in\mathcal{C}^{0}(\overline{\omega})$ and that $0\leq f^{(k)}(y)\leq 1$, for all $y\in\overline{\omega}$. Since the function $\textup{dist}(\cdot,\gamma_{0}):\overline{\omega}\to\mathbb{R}$ is Lipschitz continuous with Lipschitz constant $L=1$, we have that an application of the Rademacher theorem (cf., e.g., Theorem 1.14-2 in [10]) gives that:
$\left|\partial_{\beta}\left(\textup{dist}(\cdot,\gamma_{0})\right)\right|\leq 1,\quad\textup{ a.e. in }\omega.$
As a result, we obtain that
$\partial_{\beta}f^{(k)}(y)=\begin{cases}0&,\textup{ if }0<\textup{dist}(y,\gamma_{0})<\dfrac{1}{k},\\ \\ \leq k&,\textup{ if }\dfrac{1}{k}<\textup{dist}(y,\gamma_{0})<\dfrac{2}{k},\\ \\ 0&,\textup{ if }0\leq\textup{dist}(y,\gamma_{0})>\dfrac{2}{k},\end{cases}$
which in turn gives that $f^{(k)}\in W^{1,\infty}(\omega)$ for each integer $k\geq 1$. Since the given vector field $\bm{\eta}=(\eta_{i})$ is such that $\eta_{i}\in H^{1}(\omega)\cap L^{\infty}(\omega)$, we are in a position to apply the chain rule formula in Sobolev spaces (cf., e.g., Proposition 9.4 in [2]) so as to infer that the functions
$\eta_{i}^{(k)}:=\left(1-\dfrac{1}{k}\right)f^{(k)}\eta_{i}\in H^{1}(\omega).$
For each $s>0$, the symbol $\omega_{s}$ denotes and define the following set:
$\omega_{s}:=\{y\in\omega;\,\textup{dist}(y,\gamma_{0})>s\}.$
It is straightforward to observe that $\eta_{i}^{(k)}=0$ a.e. in $\omega\setminus\overline{\omega_{1/k}}$ and hence, a fortiori, we have that $\eta_{i}^{(k)}=0$ on $\gamma_{0}$. It is also easy to observe that, for a.a. $y\in\omega$, the geometrical constraint is valid: Indeed,
$(\bm{\theta}(y)+\eta_{i}^{(k)}(y)\bm{a}^{i}(y))\cdot\bm{q}=\dfrac{1}{k}\bm{\theta}(y)\cdot\bm{q}+\left(1-\dfrac{1}{k}\right)\left(1-f^{(k)}(y)\right)\bm{\theta}(y)\cdot\bm{q}$
$\qquad+\left(1-\dfrac{1}{k}\right)f^{(k)}(y)(\bm{\theta}(y)+\eta_{i}(y)\bm{a}^{i}(y))\cdot\bm{q}\geq\dfrac{d}{k}>0.$
Therefore, we have shown that $\bm{\eta}^{(k)}\in\bm{U}_{M}^{(2)}(\omega)$, for each integer $k\geq 1$. Let us now consider the difference
$\bm{\eta}^{(k)}-\bm{\eta}=\begin{cases}-\bm{\eta}&,\textup{ in }\omega\setminus\overline{\omega_{1/k}},\\ \\ \left[\left(1-\dfrac{1}{k}\right)f^{(k)}-1\right]\bm{\eta}&,\textup{ in }\omega_{1/k}\setminus\overline{\omega_{2/k}},\\ \\ -\dfrac{1}{k}\bm{\eta}&,\textup{ in }\omega_{2/k},\end{cases}$
and let us evaluate:
$\|\bm{\eta}-\bm{\eta}^{(k)}\|_{\bm{L}^{2}(\omega)}=\|\bm{\eta}\|_{\bm{L}^{2}(\omega\setminus\overline{\omega_{1/k}})}+\left\|\left[\left(1-\dfrac{1}{k}\right)f^{(k)}-1\right]\bm{\eta}\right\|_{\bm{L}^{2}(\omega_{1/k}\setminus\overline{\omega_{2/k}})}+\dfrac{1}{k}\|\bm{\eta}\|_{\bm{L}^{2}(\omega_{2/k})}$
$\leq\|\bm{\eta}\|_{\bm{L}^{2}(\omega\setminus\overline{\omega_{1/k}})}+2\left(\int_{\omega}|\bm{\eta}|^{2}\chi_{\omega_{1/k}\setminus\overline{\omega_{2/k}}}\,\mathrm{d}y\right)^{1/2}+\dfrac{1}{k}\|\bm{\eta}\|_{\bm{L}^{2}(\omega_{2/k})}.$
Observe that the right-hand side of the previous estimate tends to zero as $k\to\infty$; indeed, the convergence of the third term is straightforward and the convergence of the first two terms follows directly from dominated convergence. This proves that $\bm{\eta}^{(k)}\to\bm{\eta}$ in $\bm{L}^{2}(\omega)$ as $k\to\infty$.
Let us now consider the difference,
$\partial_{\beta}\eta_{i}^{(k)}-\partial_{\beta}\eta_{i}=\begin{cases}-\partial_{\beta}\eta_{i}&,\textup{ in }\omega\setminus\overline{\omega_{1/k}},\\ \\ \left[\left(1-\dfrac{1}{k}\right)f^{(k)}-1\right]\partial_{\beta}\eta_{i}+\left(1-\dfrac{1}{k}\right)(\partial_{\beta}f^{(k)})\eta_{i}&,\textup{ in }\omega_{1/k}\setminus\overline{\omega_{2/k}},\\ \\ -\dfrac{1}{k}\partial_{\beta}\eta_{i}&,\textup{ in }\omega_{2/k},\end{cases}$
and let us evaluate:
(29) $\|\partial_{\beta}\eta_{i}^{(k)}-\partial_{\beta}\eta_{i}\|_{L^{2}(\omega)}\leq\|\partial_{\beta}\eta_{i}\|_{L^{2}(\omega\setminus\overline{\omega_{1/k}})}+\dfrac{1}{k}\|\partial_{\beta}\eta_{i}\|_{L^{2}(\omega_{2/k})}$
$\qquad+\left\|\left(\left(1-\dfrac{1}{k}\right)f^{(k)}-1\right)\partial_{\beta}\eta_{i}\right\|_{L^{2}(\omega_{1/k}\setminus\overline{\omega_{2/k}})}$
$\qquad+\left\|\left(1-\dfrac{1}{k}\right)(\partial_{\beta}f^{(k)})\eta_{i}\right\|_{L^{2}(\omega_{1/k}\setminus\overline{\omega_{2/k}})}$
$\leq\|\partial_{\beta}\eta_{i}\|_{L^{2}(\omega\setminus\overline{\omega_{1/k}})}+\dfrac{1}{k}\|\partial_{\beta}\eta_{i}\|_{L^{2}(\omega_{2/k})}+2\|\partial_{\beta}\eta_{i}\|_{L^{2}(\omega_{1/k}\setminus\overline{\omega_{2/k}})}+k\|\eta_{i}\|_{L^{2}(\omega_{1/k}\setminus\overline{\omega_{2/k}})}.$
It is clear that the first three terms in the right-had side of (29) converge to zero as $k\to\infty$ by dominated convergence. The analysis of the last term is more involving and hinges on the geometrical assumptions on the set $\omega$ set forth in the statement of this theorem.
In what follows, we limit ourselves to considering case (b), i.e., the the case where $\omega$ is a rectangle and $\gamma_{0}$ is either an edge of the rectangle or contains the union of two parallel edges. For the sake of simplicity, we also assume that the origin is placed at one of the vertices of $\gamma_{0}$. We observe that, for each integer $k\geq 1$, the set $\omega\setminus\overline{\omega_{1/k}}$ is the union of rectangles such that one of their edges is one of the edges constituting the portion of the boundary $\gamma_{0}$ where the boundary conditions of place are imposed. Examples of admissible configurations for the sets $\omega$ and $\omega\setminus\overline{\omega_{2/k}}$ are depicted in Figure 3 below.
(a) The set $\gamma_{0}$ consists of one edge.
(b) The set $\gamma_{0}$ is the union of two parallel edges.
(a) The set $\gamma_{0}$ is the union of three edges.
(b) The set $\gamma_{0}$ coincides with $\gamma$.
Figure 3. The sets $\omega$ and $\omega\setminus\overline{\omega_{2/k}}$ for different instances of $\gamma_{0}$, which is here denoted by a thick black line. In (a), the set $\gamma_{0}$ coincides with one edge of the rectangular domain, in (b) the set $\gamma_{0}$ is the union of two parallel edges, in (c) the set $\gamma_{0}$ contains the union of two parallel edges and, finally, in (d), the set $\gamma_{0}$ coincides with the entire boundary $\gamma$.
The assumption on the geometry of the domain will turn out to be critical to establish the following inequality:
(30) $k\|\xi\|_{L^{2}(\omega\setminus\overline{\omega_{2/k}})}\leq C\|\xi\|_{H^{1}(\omega\setminus\overline{\omega_{2/k}})},$
for all $\xi\in H^{1}(\omega)$ such that $\xi=0$ on $\gamma_{0}$, for some $C>0$ independent of $k$. To show the inequality (30), we carry out some of the computations necessary to establish the Poincaré-Friedrichs inequality (cf., e.g., Theorem 6.5-2 in [10]). For the sake of clarity, we present the main steps for establishing the sought inequality (30). Fix $y_{0}\in\gamma_{0}$, and let $y\in\omega\setminus\overline{\omega_{2/k}}$ be such that $y=y_{0}-c\nu(y_{0})$, for some $c>0$. To fix the ideas, we consider the strip of width $2/k$ in Figure 3(a). The absolute continuity along the lines (cf., e.g., [34]) enjoyed by the function $\xi$ and the fact that $\xi(y_{0})=0$ in turn give:
$\xi(y)=\int_{y_{0,2}}^{y_{2}}\partial_{2}\xi(y_{1},t)\,\mathrm{d}t,\quad\textup{ for a.a. }y_{1}\in(0,\textup{length }\gamma_{0}).$
Consequently, we have that:
$|\xi(y)|^{2}\leq\left(\int_{y_{0,2}}^{y_{2}}|\partial_{2}\xi(y_{1},t)|\,\mathrm{d}t\right)^{2}\leq(y_{2}-y_{0,2})\int_{y_{0,2}}^{y_{2}}|\partial_{2}\xi(y_{1},t)|^{2}\,\mathrm{d}t$
$\leq(y_{2}-y_{0,2})\int_{y_{0,2}}^{y_{0,2}-\frac{2}{k}\nu_{2}(y_{0})}|\partial_{2}\xi(y_{1},t)|^{2}\,\mathrm{d}t,$
where the second inequality is due to Hölder’s inequality.
The latter estimate in turn implies that
$\int_{y_{0,2}}^{y_{0,2}-\frac{2}{k}\nu_{2}(y_{0})}|\xi(y)|^{2}\,\mathrm{d}y_{2}\leq\dfrac{2}{k^{2}}\int_{y_{0,2}}^{y_{0,2}-\frac{2}{k}\nu_{2}(y_{0})}|\partial_{2}\xi(y_{1},t)|^{2}\,\mathrm{d}t,\quad\textup{ for a.a. }y_{1}\in(0,\textup{length }\gamma_{0}),$
so that it is immediate to establish the following estimate:
$\int_{0}^{\textup{length }\gamma_{0}}\int_{y_{0,2}}^{y_{0,2}-\frac{2}{k}\nu_{2}(y_{0})}|\xi(y)|^{2}\,\mathrm{d}y_{2}\,\mathrm{d}y_{1}\leq\dfrac{2}{k^{2}}\int_{0}^{\textup{length }\gamma_{0}}\int_{y_{0,2}}^{y_{0,2}-\frac{2}{k}\nu_{2}(y_{0})}|\partial_{2}\xi(y_{1},t)|^{2}\,\mathrm{d}t$
$=\dfrac{2}{k^{2}}\|\partial_{2}\xi\|_{L^{2}(\omega\setminus\overline{\omega_{2/k}})}.$
Therefore, the claimed inequality (30) holds with $C:=\sqrt{2}$. We are now in a position to analyse the last term in the right-hand side of (29):
$k\|\eta_{i}\|_{L^{2}(\omega\setminus\overline{\omega_{2/k}})}\leq\sqrt{2}\|\eta_{i}\|_{H^{1}(\omega\setminus\overline{\omega_{2/k}})},$
and we easily observe that the right-hand side of the later term converges to zero as $k\to\infty$ by dominated convergence, thus proving that $\partial_{\beta}\eta_{i}^{(k)}\to\partial_{\beta}\eta_{i}$ in $L^{2}(\omega)$ as $k\to\infty$ and, therefore, the sought density.
The case corresponding to Figures 3(b)–(d) can be treated analogously, so as the case where the domain has smooth boundary and the boundary condition of place is imposed along the entire boundary $\gamma$.
(iii) Denote by $\mathcal{C}_{\gamma_{0}}^{\infty}(\overline{\omega})$ the space of continuously differentiable functions that vanish in a neighbourhood of $\gamma_{0}$. Then, the set $\bm{U}_{M}^{(3)}(\omega):=\bm{\mathcal{C}}_{\gamma_{0}}^{\infty}(\overline{\omega})\cap\bm{U}(\omega)$ is dense in then set $\bm{U}_{M}^{(2)}(\omega)$ with respect to the norm $\|\cdot\|_{\bm{H}^{1}(\omega)}$.
To fix the ideas, let us once again consider the case where $\omega$ is a rectangle and the boundary conditions under consideration are as in Figure 3(a), and let $\bm{\eta}\in\bm{U}_{M}^{(2)}(\omega)$ be given. Let $k>2/\delta(\bm{\eta})$ be an integer, where $\delta(\bm{\eta})$ is the positive number appearing in the definition of $\bm{U}_{M}^{(2)}(\omega)$. Let $\rho_{1/k}$ be a standard mollifier (cf., e.g., [2]) such that $\textup{supp }\rho_{1/k}\subset\subset B(0;1/k)$. Following the ideas leading to Figure 6 on page 275 of [2], we extend the original domain $\omega$ by reflection along the edges, so as to obtain a new rectangular domain $\hat{\omega}$ like the one depicted in Figure 4 below.
(a) Extension in the case where the set $\gamma_{0}$ is the union of two parallel edges.
(b) Extension in the case where the set $\gamma_{0}$ consists of one edge.
(c) Extension in the case where the set $\gamma_{0}$ consists of one edge. Reflection in proximity of a corner of $\omega$.
Figure 4. The set $\hat{\omega}$ is an extension by reflection of the original domain $\omega$. First, we reflect the strips of width $\delta(\bm{\eta})$ across the edge of $\gamma_{0}$ it is adjacent to. Second, we reflect the strips of width $\delta(\bm{\eta})$ across the edges of the intermediary extension that contain the portion of boundary $\gamma\setminus\gamma_{0}$. The hatched patterns denote the portion of $\omega$ where the given function $\bm{\eta}\in\bm{U}_{M}^{(2)}(\omega)$ vanishes. The dotted pattern denotes the points in the domain extension where the extension of $\bm{\eta}$ by reflection vanishes (cf., e.g., Theorem 9.7 in [2]). The point $\hat{y}$, which is represented by a black dot, is the antipodal point corresponding to $y\in\omega$ with respect to $y_{0}\in\gamma$, denoted by a hallow dot. This means that there exists a number $c>0$ such that $\hat{y}=y+2c\nu(y_{0})$. In the case illustrated in (a), where the set $\gamma_{0}$ is the union of two parallel edges, the extension vanishes near the corners of $\omega$ and the extension by reflection of the contravariant basis with respect to the remaining edges is performed in the manner described at the beginning of section 6. If the set $\gamma_{0}$ consists of one edge only, then the extension by reflection of the contravariant basis is performed in the usual manner when $y_{0}$ is not a corner point (b). Otherwise, the extension by reflection of the contravariant basis is carried out with respect to two points $y_{1}$ and $\hat{y}_{2}$ that are antipodal with respect to the corner point $y_{0}$ under consideration (c). The latter is equivalent to considering the extension by reflection of the contravariant basis at the point $y_{1}$ to the point $\hat{y}_{1}$ first, and then considering the extension by reflection of the contravariant basis at the point $\hat{y}_{1}$ to the point $\hat{y}_{2}$.
After multiplying the extension by reflection of $\bm{\eta}$ to $\hat{\omega}$ (cf., e.g., Theorem 9.7 of [2]) by an appropriate cut-off function which is equal to 1 in a set $\omega_{1}$ such that $\omega\subset\subset\omega_{1}\subset\subset\hat{\omega}$, we infer that the regularisation of $\eta_{i}$, which are denoted and defined by
$\eta_{i}^{(k)}:=\rho_{1/k}\ast\eta_{i},$
are of class $\mathcal{C}^{\infty}(\overline{\omega})$ and satisfy $\eta_{i}^{(k)}=0$ in $\omega\setminus\overline{\omega_{\delta(\bm{\eta})/2}}$. Moreover, the classical regularisation theory (cf., e.g., [2]) gives that:
$\eta_{i}^{(k)}\to\eta_{i},\quad\textup{ in }H^{1}(\omega)\textup{ as }k\to\infty.$
To complete the proof, we need to verify that $\bm{\eta}^{(k)}$ satisfies the geometrical constraint ensuring that the deformed reference configuration of the shell is contained in the prescribed half-space. The case where $\textup{dist}(y,\gamma)>\delta(\bm{\eta})/2$ is straightforward to treat, as the support of the mollifier $\rho_{1/k}$ does not invade the complement of $\omega$. To complete the proof, let us then consider the case where $y\in\omega$ such that $\textup{dist}(y,\gamma)\leq\delta(\bm{\eta})/2$ and, in the same spirit as Figure 4, let us consider the point $\hat{y}$ that is specular to $y$ with respect to the edge of $\gamma$ whose distance from $y$ is minimal.
The definition of $\bm{U}_{M}^{(2)}(\omega)$, the assumed properties for $\hat{\bm{\theta}}$, assumption (17), the property of the reflection and the fact that $\eta_{i}$ extends by zero in the nearness of the edges where the boundary condition of place is imposed give:
(31) $\left(\bm{\theta}(y)+\eta_{i}^{(k)}(y)\bm{a}^{i}(y)\right)\cdot\bm{q}=\bm{\theta}(y)\cdot\bm{q}+\left(\int_{B(0;1/k)}\rho_{1/k}(z)\hat{\eta}_{i}(y-z)\,\mathrm{d}z\right)\bm{a}^{i}(y)\cdot\bm{q}$
$=\int_{B(0;1/k)}\rho_{1/k}(z)\left(\hat{\bm{\theta}}(y-z)+\hat{\eta}_{i}(y-z)\hat{\bm{a}}^{i}(y-z)\right)\cdot\bm{q}\,\mathrm{d}z$
$\quad+\int_{B(0;1/k)}\rho_{1/k}(z)(\hat{\bm{\theta}}(y)-\hat{\bm{\theta}}(y-z))\cdot\bm{q}\,\mathrm{d}z+\int_{B(0;1/k)}\rho_{1/k}(z)\hat{\eta}_{i}(y-z)(\hat{\bm{a}}^{i}(y)-\hat{\bm{a}}^{i}(y-z))\cdot\bm{q}\,\mathrm{d}z$
$\geq d\delta(\bm{\eta})-\max_{z\in B(0;1/k)}|\hat{\bm{\theta}}(y)-\hat{\bm{\theta}}(y-z)|-\|\hat{\eta}_{i}\|_{L^{2}(\mathbb{R}^{2})}\left(\max_{\begin{subarray}{c}1\leq i\leq 3\\ z\in B(0;1/k)\end{subarray}}\left|\hat{\bm{a}}^{i}(y)\cdot\bm{q}-\hat{\bm{a}}^{i}(y-z)\cdot\bm{q}\right|\right).$
The continuity of $\hat{\bm{\theta}}$ and the continuity property established in (17) in turn give that the second and third addend in the right-hand side of (31) tend to zero as $k\to\infty$. As a result, for $k$ sufficiently large, we obtain that:
$\left(\bm{\theta}(y)+\eta_{i}^{(k)}\bm{a}^{i}(y)\right)\cdot\bm{q}\geq 0,$
so that $\bm{\eta}^{(}k)\in\bm{U}_{M}^{(3)}(\omega)$ by the arbitrariness of $y\in\omega$.
(iv) Conclusion of the proof.
Upon completion of parts (i)-(iii), we have shown that, for a given $\bm{\eta}\in\bm{U}(\omega)$, there exists a sequence $\left\{\bm{\eta}^{(k)}\right\}_{k\geq 1}\subset\bm{\mathcal{C}}^{\infty}_{\gamma_{0}}(\overline{\omega})\cap\bm{U}(\omega)$ such that:
$\bm{\eta}^{(k)}\to\bm{\eta},\quad\textup{ in }\bm{H}^{1}(\omega)\textup{ as }k\to\infty.$
Thanks to the inequality (18), we have that $\left|\bm{\eta}^{(k)}-\bm{\eta}\right|_{\omega}^{M}\to 0$, as $k\to\infty$, thus showing that:
$\overline{\bm{\mathcal{C}}^{\infty}_{\gamma_{0}}(\overline{\omega})\cap\bm{U}(\omega)}^{|\cdot|_{\omega}^{M}}=\bm{U}(\omega).$
Therefore, the claim $\bm{U}_{M}^{\sharp}(\omega)=\bm{U}^{\sharp}(\omega)$ immediately follows by observing that the iterates constructed in (iii) vanish near $\gamma_{0}$, thus satisfying a homogeneous boundary condition of Neumann type along $\gamma_{0}$. ∎
We now comment the assumption of geometrical nature set forth in the statement of Theorem 6.3.
To begin with, we comment the assumption on the shape of the domain $\omega$ and of the portion $\gamma_{0}$ of the boundary $\gamma$ where the boundary conditions of place are imposed. Observe that if $\gamma$ is smooth and $\textup{meas}(\gamma\setminus\gamma_{0})>0$, then the set $\omega\setminus\overline{\omega_{2/k}}$ considered in step (ii) of the proof would include a portion of the set $\gamma\setminus\gamma_{0}$. By so doing, it is not possible to carry out the calculations leading to the inequality (30) since the function under consideration does not vanish on $\gamma\setminus\gamma_{0}$.
Second, assuming that $\omega$ is a rectangle and $\gamma_{0}$ is the union of entire edges allows us to apply the argument based on the absolute continuity along the lines for establishing part (ii) of Theorem 6.3. For general polygons and for instances of $\gamma_{0}$ that cover only a portion of an edge, these computations cannot be carried out.
Third, and finally, the fact that the linearly elastic generalised membrane shells we are considering are of the first kind renders the mapping $|\cdot|_{\omega}^{M}$ a norm on $\bm{V}(\omega)$, thus ensuring the claimed density.
A first example of middle surface satisfying the assumptions critical to establish Theorem 6.3 is, for instance, a portion of a spherical cap such that the portion of contour where the boundary conditions of place are imposed lies on a plane that is parallel to the plane associated with the orthogonal complement of the given unit-vector $\bm{q}$. For $\bm{q}=(0,0,1)$, a parametrisation for one such surface is given by
$y=(y_{1},y_{2})\in[0,\pi]\times\left[0.1,c\dfrac{\pi}{2}\right]\mapsto\bm{\theta}(y):=\left(\cos y_{1}\sin y_{2},\sin y_{1}\sin y_{2},\cos y_{2}\right),\quad\textup{ where }0<c<1,$
and the portion $\gamma_{0}$ of the boundary $\gamma$ where the boundary conditions are imposed is given by $\gamma_{0}:=[0,\pi]\times\{0.1\}$.
A second example of middle surface satisfying the assumptions critical to establish Theorem 6.3 is, for instance, a portion of a cylinder whose straight boundary edges lie on the same plane, and this plane these edges lie on is parallel to plane associated with the orthogonal complement of the given unit-vector $\bm{q}$. For $\bm{q}=(0,0,1)$, a parametrisation for one such surface is given by
$y=(y_{1},y_{2})\in[0.1,\pi-0.1]\times[0,2]\mapsto\bm{\theta}(y):=\left(\cos y_{1},y_{2},\sin y_{1}\right),$
and the portion $\gamma_{0}$ of the boundary $\gamma$ where the boundary conditions are imposed is given by $\gamma_{0}:=[0.1,\pi-0.1]\times\{-2\}$.
A third example of middle surface satisfying the assumptions critical to establish Theorem 6.3 is, for instance, a portion of a one-sheeted hyperboloid of revolution whose boundary is the union of two circles, and these circles lie on two distinct planes which are both either orthogonal to the given unit-vector $\bm{q}$ or they are orthogonal to the orthogonal complement of the unit-vector $\bm{q}$. For $\bm{q}=(0,1,0)$, a parametrisation for one such surface is given by
$y=(y_{1},y_{2})\in[0.1,\pi-0.1]\times[-2,2]\mapsto\bm{\theta}(y):=\left(\sqrt{1+y_{2}^{2}}\cos y_{1},\sqrt{1+y_{2}^{2}}\sin y_{1},y_{2}\right),$
and the portion $\gamma_{0}$ of the boundary $\gamma$ where the boundary conditions are imposed is given by $\gamma_{0}:=(\{0\}\times[-2,2])\cup(\{\pi-0.1\}\times[-2,2])$.
Despite the assumptions of geometrical nature required to bring the proof of Theorem 6.3 to a completion, the surfaces satisfying these assumptions are non-trivial in number and in type.
Figure 5. Three examples of surfaces defined over a rectangular domain $\omega\subset\mathbb{R}^{2}$ to which Theorem 6.3 can be applied. (a) A portion of a spherical cap such that the contour where the boundary conditions of place are imposed lies on a plane parallel to the orthogonal complement of $\bm{q}=(0,0,1)$; (b) A portion of a cylinder such that the image of the edges of $\gamma$ that are orthogonal to $\gamma_{0}$ lies on a plane that is parallel to the orthogonal complement of $\bm{q}=(0,0,1)$; (c) A portion of a hyperboloid of revolution such that the symmetry axis director is parallel to $\bm{q}=(0,1,0)$. In order to apply Theorem 5.1, we need to require the validity of condition (3). Note that this is in line with the conclusion obtained for linearly elastic elliptic membrane shells [20]. Assuming the validity of (3) certainly affects the choices for the middle surfaces for which the asymptotic analysis presented in Theorem 5.1 holds. The density result presented in Theorem 6.3 holds, in general, for a larger class of surfaces.
Conclusion and final remarks
In this paper we identified a set of two-dimensional variational inequalities that model the displacement of a linearly elastic generalised membrane shell of the first kind subjected to a confinement condition, expressing that all the points of the admissible deformed configurations remain confined in a prescribed half-space.
The starting point of the rigorous asymptotic analysis we carried out is a set of variational inequalities based on the classical energy of three-dimensional linearised elasticity, and posed over a non-empty, closed and convex subset of a suitable Sobolev space. A rigorous asymptotic analysis departing from one such three-dimensional model led to the individuation of a two-dimensional model posed over an abstract completion of a non-empty, closed and convex subset of the space $\bm{H}^{1}(\omega)$. We observed the two-dimensional variational inequalities recovered as a result of the aforementioned asymptotic analysis are the same as the one recovered in [17, 16], where the departure point was Koiter’s model subjected to the same confinement condition considered in this paper.
However, the sets where the solutions for the two-dimensional limit model are sought appear not to be the same, in the sense that the asymptotic analysis departing from the three-dimensional model with obstacle and the asymptotic analysis departing from Koiter’s model with obstacle individuate two sets denoted by $\bm{U}_{M}^{\sharp}(\omega)$ and $\bm{U}^{\sharp}(\omega)$, respectively, which in general might not coincide.
The result announced in Theorem 6.3 establishes that under suitable assumptions on the geometry of the linearly elastic generalised membrane shell of the first kind under consideration, these sets actually do coincide, and we are thus in a position to regard the justification of Koiter’s model for linearly elastic generalised membrane shells of the first kind subjected to remaining confined in a prescribed half-space as complete.
Finally, in connection with the results presented in [16, 17, 19, 20], we observe that if the shell middle surface is elliptic and $\gamma_{0}=\gamma$, then the linearly elastic shell under consideration is a linearly elastic elliptic membrane shell so that, thanks to Korn’s inequality for elliptic surfaces (cf., e.g., Theorem 2.7-3 in [8]), we have that $\bm{U}_{M}^{\sharp}(\omega)=\bm{U}_{M}(\omega)=\{\bm{\eta}=(\eta_{i})\in H^{1}_{0}(\omega)\times H^{1}_{0}(\omega)\times L^{2}(\omega);(\bm{\theta}+\eta_{i}\bm{a}^{i})\cdot\bm{q}\geq 0\textup{ a.e. in }\omega\}$, provided that $\min_{y\in\overline{\omega}}(\bm{\theta}(y)\cdot\bm{q})>0$ and that $\min_{y\in\overline{\omega}}(\bm{a}^{3}(y)\cdot\bm{q})>0$. Note that, differently from the case of linearly elastic elliptic membrane shells, the condition $\min_{y\in\overline{\omega}}(\bm{a}^{3}(y)\cdot\bm{q})>0$, here denoted by (3), has to be assumed for being able to carry out the asymptotic analysis presented in Theorem 5.1 while, in [19, 20], this condition has to be assumed to establish the validity of the “density property”.
Declarations
Authors’ Contribution
Not applicable.
Acknowledgements
Not applicable.
Ethical Approval
Not applicable.
Availability of Supporting Data
Not applicable.
Competing Interests
All authors certify that they have no affiliations with or involvement in any organisation or entity with any competing interests in the subject matter or materials discussed in this manuscript.
Funding
This research was partly supported by the University Development Fund of The Chinese University of Hong Kong, Shenzhen.